### v1

# Noise-tolerant public-key quantum money from a classical oracle

Abstract
Quantum money is the task of verifying the validity of banknotes while ensuring that they cannot be counterfeited. Public-key quantum money allows anyone to perform verification, while the private-key setting restricts the ability to verify to banks, as in Wiesner’s original scheme. The current state of technological progress means that errors are impossible to entirely suppress, hence the requirement for noise-tolerant schemes. We show for the first time how to achieve noise-tolerance in the public-key setting. Our techniques follow Aaronson and Christiano’s oracle model, where we use the ideas of quantum error correction to extend their scheme: a valid banknote is now a subspace state possibly affected by noise, and verification is performed by using classical oracles to check for membership in “larger spaces.” Additionally, a banknote in our scheme is minted by preparing conjugate coding states and applying a unitary that permutes the standard basis vectors.
1 Introduction
The history of quantum money is essentially as long as quantum information itself. At its core, quantum money aims to address the issue of verifying the validity of banknotes while ensuring that they cannot be counterfeited (i.e., cloned). Famously, this problem was first addressed by Wiesner in [Wie83] where, put simply, a single banknote consisted of $n$ conjugate coding states, and verification involved measuring each qubit in the basis that it was prepared in (thus, this verification procedure requires the verifier to know how a given banknote was prepared). Without knowledge of how the banknote was prepared, a counterfeiter attempting to produce a second copy of a valid banknote has a success probability of at most $(3/4)^{n}$, as shown in [MVW13]. Informally, Wiesner’s scheme is secure within a model where the verification procedure only says whether the banknote is accepted or rejected (the post-measurement state is not returned).
Crucially, in Wiesner’s scheme, only banks are capable of verifying banknotes (since the ability to verify enables the minting of new, valid banknotes). Quantum money where anyone can perform verification of a banknote is known as public-key quantum money. Public-key quantum money was first addressed in [Aar09] where it was proved that there exists a quantum oracle relative to which secure public-key quantum money is possible. Then, in the foundational work of [AC12], it was proven that public-key quantum money relative to a classical oracle is secure. The scheme involves a banknote as a subspace state |$\ket{A}$ (a uniform superposition of the vectors in a random $n/2$-dimensional subspace $A\subseteq\mathbb{F}^{n}_{\_}2$) and verification is performed through classical oracles that check for membership in $A$ and the dual subspace $A^{\perp}$. The technical result from [AC12] is that any quantum algorithm that maps |$\ket{A}$ to |$\ket{A}^{\otimes 2}$ must make exponentially many queries, implying that a quantum polynomial-time adversary with access to the verification procedure would not be able to counterfeit a banknote.
Noise-tolerant schemes:
While much work has been devoted to public-key quantum money, that work has been focused on instantiating the oracles of [AC12] or devising entirely new schemes. Regardless of how public-key quantum money is obtained in the plain model (without oracles), it will be necessary for such a scheme to be noise-tolerant (i.e., robust) for it to be practically meaningful. Indeed, one of the most pressing practical issues of quantum money is the necessity of quantum storage for the banknote (for a review of quantum storage, see [HEH+16]; for more recent work, see [WLQ+21] where the authors maintain coherence of a qubit for over one hour). Due to the inherently sensitive nature of quantum information, any amount of time in quantum storage almost guarantees the incursion of quantum errors. If noise affects an honest banknote to the point that it does not pass verification, then it has lost its prescribed value. In the classical oracle scheme of [AC12], a single error will cause an honest banknote to no longer pass verification (see section 3.2). To the best of our knowledge, we are unaware of any previous work that focuses on noise-tolerant public-key quantum money.
In the private-key setting of quantum money, there has indeed been work towards noise-tolerance (see [PYJ+12, AA17, Kum19]). Notably, the scheme in [AA17] is based off hidden matching quantum retrieval games; in this model, the authors prove that, for $n=14$, the maximum achievable noise-tolerance is $23.3\%$, which makes their protocol nearly optimal since it can tolerate up to $23.03\%$ noise. In [BSS21], the authors construct a noise-tolerant private tokenized signature scheme (from which one can obtain private-key quantum money; see section 1.3). To simply describe their approach, a token is a tensor product of conjugate coding states, signing the bit 00 ($1$) involves measuring all qubits in the computational (Hadamard) basis, and verification amounts to comparing the measured string with a bank’s secret string to check for agreement; noise-tolerance comes from allowing inconsistencies in a fraction of the string. Unfortunately, their scheme is not secure against an adversary that can query the classical verification oracle in quantum superposition.
1.1 Contributions
We present a scheme for noise-tolerant public-key quantum money and analyze its completeness and soundness error. Intuitively, completeness error quantifies the correctness of the scheme while soundness error quantifies the success probability of a counterfeiter. More precisely, completeness error $\varepsilon$ means that valid banknotes (those minted by the bank and then possibly affected by tolerable noise) are accepted by the verification procedure with probability at least $1-\varepsilon$. Soundness error $\delta$ means that a counterfeiter, who possesses a valid banknote and whose quantum circuit is polynomial in size, can produce another banknote such that both banknotes are accepted with probability at most $\delta$. We show that our scheme has perfect completness error and has |_$|\mathcal{E}_{\_}q|^{2}/\textsf{exp}(n)$ soundness error, where _$\mathcal{E}_{\_}q$ represents the set of tolerated Pauli $X$ and $Z$ errors, each affecting up to $q$ of the $n$ qubits in our banknote (i.e., this corresponds to tolerating $q$ arbitrary errors). A freshly minted banknote in our scheme is a subspace state, as in [AC12], though we use a relation formalized in [CV22] to show that such a banknote can be prepared by generating conjugate coding states and applying a unitary that permutes the standard basis vectors. Noise-tolerance is built into our verification procedure which uses classical oracles to check for membership in spaces that are determined by $q$ and the chosen subspace.
1.2 Techniques
Our approach to noise-tolerance starts with the fact that the banknote state used in [AC12] can be viewed as a codeword for a quantum error correcting code; specifically, a Calderbank-Shor-Steane (CSS) code. With this connection, we ask: is it possible to securely use the corresponding CSS code to detect and tolerate errors? Ignoring security, momentarily, the answer is yes. In the stabilizer formalism [Got97], error detection involves measuring a set of stabilizer generators; the resulting classical string, known as a syndrome, can be processed to tell us what errors occurred and on which qubits. This information can be used to correct the noise or, more simply, to accept a state that has been afflicted by correctable noise. When it comes to security, we find that direct application of this error detection procedure, without “hiding” its properties, can compromise security, as the generators reveal information about the underlying subspace (see section 3.3.1).
The next and perhaps most obvious approach is to apply a layer of error correction on top of the oracle scheme of [AC12]. With this approach, one would encode the subspace state |$\ket{A}$ and then proceed in one of two ways: perform verification in a fault-tolerant manner; or, error correct, decode, and then apply the normal verification procedure. In section 3.3.2, we briefly discuss this approach and outline some of its obstacles, though we do not rule it out as a possibility and instead leave it as one of the open questions listed in section 5.
In this paper, the direction we take is to design a scheme that is inherently noise-tolerant (although we focus on tolerating noise, we discuss in remark 4.10 how our scheme could be easily adapted to also correct the errors). We present two approaches for a noise-tolerant scheme, both of which achieve the same completeness and soundness error. While the two approaches offer similar results, we include both of them in this paper since we cannot definitively say which would be preferable in the plain model (see section 5 where we discuss this as a direction for future research). Notably, both approaches can be seen as natural extensions of the original classical oracle scheme of [AC12]. The difference between the two approaches, described simply, is that one approach checks for membership in a union of subspaces while the other approach checks membership in a union of subsets. Note that in both approaches, we tolerate up to $q$ arbitrary errors by tolerating up to $q$ Pauli-$X$ errors and up to $q$ Pauli-$Z$ errors, as is done in CSS codes.
Conceptually, the subspace approach closely resembles error detection in the stabilizer formalism. It starts by recognizing that each tolerable syndrome (i.e., tolerable error) corresponds to a subspace. We can then form two unions of disjoint subspaces (one for Pauli-$X$ errors and one for Pauli-$Z$ errors). We then perform verification by using classical oracles to check if the banknote is a member of these two unions.
In the subset approach, we do essentially the same thing except that we check membership in subsets. We recognize that, in the computational basis, a Pauli-$X$ error causes a subspace state to become a superposition of the vectors in a coset (i.e., it becomes a coset state), where each Pauli-$X$ error corresponds to a unique coset. Under this relation, the set of tolerated Pauli-$X$ errors corresponds to a collection of cosets. To then accept a state corrupted by a Pauli-$X$ error, we use a classical oracle to check membership in the subset defined as the union of all cosets from this collection. We tolerate Pauli-$Z$ errors by checking membership in a similarly defined subset.
We focus primarily on the subset approach, as it is conceptually simpler and the security proof for the subspace approach is almost the same (throughout the security analysis, we remark on any changes that would be necessary for the subspace approach). As one might expect, for a banknote with a fixed size $n$, the more noise we tolerate the easier it will be for a counterfeiter to succeed; this is captured by the |_$|\mathcal{E}_{\_}q|^{2}/\textsf{exp}(n)$ soundness error.
On top of the above construction, we make use of an observation (formalized in [CV22]) on the relation between conjugate coding states and coset states in the generation of our banknotes. Since conjugate coding states are practically simple to prepare, we use this relation with the intent of making banknote generation similarly simple. This relation (2.1), allows the bank to generate the subspace state banknote by preparing conjugate coding states and then applying a unitary that permutes the standard basis vectors. Intuitively, this does not affect the security of our scheme, as an adversary still only receives a subspace state from the bank.
1.3 Further background on quantum money
As noted earlier, Wiesner’s scheme is secure in a model where the post-measurement state is not returned. In a model where the verification procedure returns accept/reject and the post-measurement state, there is a simple attack on Wiesner’s scheme where, by repeatedly submitting banknotes for verification, a counterfeiter can produce a second banknote [Lut10, Aar09]. More generally, quantum state restoration [FGH+10] can be used to break Wiesner’s scheme and schemes like it.
There have been several attempts to obtain public-key quantum money in the plain model, though this has proven to be a highly non-trivial task. Most attempts at instantiating the oracles have subsequently been broken. In the quantum oracle model of [Aar09], Aaronson proposed an explicit scheme based on random stabilizer states, though this was later broken in [LAF+10]. When the authors of [AC12] introduced the classical oracle model, they also proposed an instantiation that was later broken in [CPFP15, CPDDF+19, BDG23]. In [Zha19], Zhandry proposed a new primitive called quantum lightning from which one can obtain public-key quantum money; informally, quantum lightning is like public-key quantum money except that even banknote generation is public, though an adversary is still unable to produce two banknotes with the same serial number. Zhandry also gave a construction of quantum lightning in [Zha19], though, as noted in a later version of the paper [Zha21], a weakness in the underlying hardness assumption was detected in [Rob21]. A full attack on the scheme was eventually found in [BDG23]. Also in [Zha19], Zhandry instantiated the oracles in [AC12] by using indistiguishability obfuscation (iO) [BGI+12, GGH+13, SW14]. The existence of post-quantum iO remains unsolved, as candidate constructions proposed in [WW21, GP21] have been subject to attacks [HJL21], but results such as [BGMZ18] suggest that it could still exist.
Apart from the previously mentioned schemes that focus on instantiating the oracles, there is a category of proposed public-key quantum money schemes that rely on mostly untested hardness assumptions. A particularly well-known example is a scheme from [FGH+12] that relies on knot theory; while the scheme remains unbroken, it is difficult to analyze and lacks a security proof. Also in this category are the schemes presented in [Kan18, KSS21], though some analysis on [KSS21] has been done in [BDG23].
In [LMZ23, AHY23], the authors push us towards a better understanding of suitable problems for instantiating public-key quantum money: in [LMZ23], the authors rule out a natural class of quantum money schemes that rely on lattice problems; and in [AHY23], the authors rule out a class of money schemes that are constructed from black box access to collision-resistant hash functions (where the verification procedure only makes classical queries to the hash function).
It is also worth mentioning the notion of quantum tokenized signature schemes. The relationship of this primitive to quantum money is characterized in [BDS23] where the authors show that a testable public (private) tokenized signature scheme can give public (private) quantum money. The testable tokenized signature scheme constructed in [BDS23] is based on the oracle scheme in [AC12]. In [CLLZ21], the authors construct a variant of the [BDS23] scheme (coset states are used instead of subspace states) and show its security in the plain model by using post-quantum iO. The schemes in [BDS23] and [CLLZ21] are not noise-tolerant, and we are unaware of any noise-tolerant public tokenized signature scheme.
1.4 Organization
The paper is organized in the following manner. In section 2, we review relevant material on quantum information, classical oracles, and quantum error correction. In section 3, we review and introduce definitions for quantum money, and we specify our noise-tolerant quantum money scheme. In section 4, we present the security analysis of our scheme. In section 5, we list a few open questions for future research.
2 Preliminaries
We say $f(n)\in\Omega(g(n))$ if there exist constants $c,n_{\_}0$ such that
$cg(n)\leq f(n),\quad\forall n>n_{\_}0.$
Given two bit strings $x,y\in\{0,1\}^{n}$, the notation $x\|y$ denotes the concatenation of the two bit strings, i.e., $x\|y\coloneqq(x,y)\in\{0,1\}^{2n}$.
For a subspace $A$, we denote its dual (i.e., orthogonal complement) by $A^{\perp}$.
2.1 Quantum information
Let $x\in\{0,1\}^{n}$ and let $\mathsf{H}$ denote the Hadamard matrix. The $n$-qubit Hadamard transfom is given by
$\mathsf{H}^{\otimes n}\ket{x}=\frac{1}{\sqrt{2^{n}}}\sum_{\_}{z\in\{0,1\}^{n}}(-1)^{x\cdot z}\ket{z}$
An arbitrary finite-dimensional Hilbert space is denoted as $\mathcal{H}$. A density operator $\rho$ on $\mathcal{H}$ is a positive semidefinite operator on $\mathcal{H}$ with trace one, $\operatorname{Tr}(\rho)=1$.
The single-qubit Pauli-$X$ and -$Z$ operators are, respectively, given by
$X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\quad\text{ and }\quad Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$
For a linear subspace $A$ of $\mathbb{F}^{n}_{\_}2$ and $t,t^{\prime}\in\{0,1\}^{n}$, a subspace state |$\ket{A}$ and a coset state |$\ket{A_{\_}{t,t^{\prime}}}$ are defined as
|$\ket{A}=\frac{1}{\sqrt{|A|}}\sum_{\_}{v\in A}\ket{v}\quad\text{and}\quad\ket{A_{\_}{t,t^{\prime}}}=X^{t}Z^{t^{\prime}}\ket{A}=\frac{1}{\sqrt{|A|}}\sum_{\_}{v\in A}(-1)^{v\cdot t^{\prime}}\ket{v+t},$ (1)
where $X^{t}=X^{t_{\_}1}\otimes\cdots\otimes X^{t_{\_}n}$, $Z^{t^{\prime}}=Z^{t^{\prime}_{\_}1}\otimes\cdots\otimes Z^{t^{\prime}_{\_}n}$. The subspace state is a superposition over all the elements in the subspace $A$. In the computational basis, the coset state is a superposition over all elements in the coset $A+t$, while in the Hadamard basis, it is a superposition over all elements in the coset $A^{\perp}+t^{\prime}$.
For $x,\theta\in\{0,1\}^{n}$, an $n$-tensor product of conjugate coding states is
|$\ket{x}_{\_}{\theta}=\ket{x_{\_}1}_{\_}{\theta_{\_}1}\cdots\ket{x_{\_}n}_{\_}{\theta_{\_}n},$ (2)
where |$\ket{x_{\_}i}_{\_}{\theta_{\_}i}=\mathsf{H}^{\theta_{\_}i}\ket{x_{\_}i}$ and $\mathsf{H}$ is the Hadamard gate.
The following lemma formally establishes the relationship between coset states and conjugate coding states. For this lemma, let =$\mathcal{B}=\{u_{\_}1,\ldots,u_{\_}n\}$ be a basis of $\mathbb{F}^{n}_{\_}2$ and let $U_{\_}\mathcal{B}$ be a unitary of $(\mathbb{C}^{2})^{\otimes n}$ which permutes the standard basis vectors:
$\forall x\in\{0,1\}^{n},\quad\quad U_{\_}\mathcal{B}\ket{x}=\Ket{\sum_{\_}ix_{\_}iu_{\_}i}.$ (3)
Lemma 2.1 ([CV22]).
Let $\{u_{\_}1,\ldots,u_{\_}n\}$ be a basis of $\mathbb{F}^{n}_{\_}2$ and $U_{\_}\mathcal{B}$ be defined as in equation 3. Let $\{u^{1},\ldots,u^{n}\}$ be its dual basis, i.e., $u^{i}\cdot u_{\_}j=\delta_{\_}{i,j}$. Let $T\subseteq\{1,\ldots,n\}$ be such that $|T|=n/2$ and $A=\text{Span}\{u_{\_}i:i\in\overline{T}\}$. Let $\theta\in\{0,1\}^{n}$ be the indicator of $\overline{T}$ (i.e., $\theta_{\_}i=1$ if and only if $i\notin T$), and $x\in\{0,1\}^{n}$. Let $t=\sum_{\_}{i\in T}x_{\_}iu_{\_}i$ and $t^{\prime}=\sum_{\_}{i\in\overline{T}}x_{\_}iu^{i}$. Then
|$\ket{A_{\_}{t,t^{\prime}}}=U_{\_}\mathcal{B}\ket{x}_{\_}\theta.$ (4)
The fidelity of quantum states $\rho$ and $\sigma$ is defined as
$F(\rho,\sigma)\coloneqq\operatorname{Tr}{\sqrt{\rho^{1/2}\sigma\rho^{1/2}}}.$ (5)
When we take the fidelity of $\rho$ and a pure state |$\ket{\psi}$, equation 5 reduces to
$F(\rho,\ket{\psi})=\sqrt{\braket{\psi}{\rho}{\psi}}.$
Furthermore, by Uhlmann’s theorem,
$F(\rho,\sigma)=\max_{\_}{\ket{\psi},\ket{\phi}}\left|\braket{\psi}{\phi}\right|,$
where the maximization is over all purifications |$\ket{\psi},\ket{\phi}$ of $\rho,\sigma$, respectively. The fidelity of $\rho$ with a subspace $S$ is defined as
$F(\rho,S)=\max_{\_}{\ket{\psi},\ket{\phi}}\left|\braket{\psi}{\phi}\right|$ | (6)
where the maximization is over all purifications |$\ket{\psi}$ of $\rho$ and all unit vectors |$\ket{\phi}\in S$.
Lemma 2.2 (“Triangle inequality” for fidelity [AC12]).
Suppose $\braket{\psi}{\rho}{\psi}\geq 1-\epsilon$ and $\braket{\phi}{\sigma}{\phi}\geq 1-\epsilon$. Then $F(\rho,\sigma)\leq|\braket{\psi}{\phi}|+2\epsilon^{1/4}$.
2.2 Classical oracles
Let $f:\{0,1\}^{*}\rightarrow\{0,1\}$ be some Boolean function. Then a classical oracle $U$ is a unitary transformation of the form
$U\ket{x}\bigg{(}\frac{\ket{0}-\ket{1}}{\sqrt{2}}\bigg{)}$ =|$=\ket{x}\bigg{(}\frac{\ket{0\oplus f(x)}-\ket{1\oplus f(x)}}{\sqrt{2}}\bigg{)}$
$\Leftrightarrow U\ket{x}\ket{-}$ $=(-1)^{f(x)}\ket{x}\ket{-}.$
Since the ancillary qubit is left untouched by this operation, we refrain from writing it and simply say $U\ket{x}=(-1)^{f(x)}\ket{x}$. Frequently, we use a classical oracle $U_{\_}A$ as a means for checking membership in some subset $A$ in the following sense
$U_{\_}A\ket{x}=\begin{cases}-\ket{x},\quad\text{if }x\in A\\ \ket{x},\quad\text{otherwise}.\end{cases}$
As noted in [AC12], we can implement $U_{\_}A$ as a projector _$\mathbb{P}_{\_}A$. To do this, we initialize a control qubit $\ket{+}$ to control the application of $U_{\_}A$ to |$\ket{x}$. After this, we measure the control qubit in the Hadamard basis. If the outcome is $\ket{-}$, then $x\in A$, otherwise $x\notin A$. Indeed,
[…]
Note that, unless otherwise stated, a classical oracle can be queried in a quantum superposition.
2.3 Quantum error correction and CSS codes
In this section, we follow [NC10] to review a well-known category of quantum error correcting codes known as Calderbank-Shor-Steane (CSS) codes. Before proceeding, it is worth pointing out that we focus on correcting bit-flip (Pauli-$X$) errors and phase-flip (Pauli-$Z$) errors. This is sufficient because, by the discretization of quantum errors, the ability to correct bit- and phase-flip errors means that an arbitrary quantum error can be corrected.
CSS codes can be constructed from classical linear codes. An $[n,k]$ classical linear code $C$ that encodes $k$ bits into an $n$-bit codespace is specified by an $n\times k$ generator matrix $G$ whose entries are elements of _$\mathbb{Z}_{\_}2=\{0,1\}$. The columns of $G$ are linearly independent and span the codespace.
Equivalently, an $[n,k]$ code can be defined as the kernel of an $(n-k)\times n$ matrix $H$ whose entries are elements of _$\mathbb{Z}_{\_}2$. That is, the codespace consists of all $x\in\{0,1\}^{n}$ such that $Hx=0$. The matrix $H$ is known as a parity check matrix.
The distance of a code $C$ is defined as
$d(C)\coloneqq\min_{\_}{x,y\in C,x\neq y}d(x,y).$
where $d(x,y)$ is the Hamming distance. Since $d(x,y)=\text{wt}(x+y)$, where $\text{wt}(\cdot)$ is the Hamming weight, we equivalently have
$d(C)=\min_{\_}{x\in C,x\neq 0}\text{wt}(x).$ (7)
If a code has distance $d\geq 2q+1$ for some integer $q$, then the code is able to correct up to $q$ errors (for a noisy codeword $y^{\prime}$, there is a unique codeword $y$ that is within Hamming distance $q$ of $y^{\prime}$).
Given an $[n,k]$ classical linear code $C$ with generator matrix $G$ and parity check matrix $H$, one can define the dual of $C$, denoted as $C^{\perp}$, as the code with generator matrix $H^{T}$ and parity check matrix $G^{T}$.
For quantum codes, we use double square brackets $[[n,k]]$ to differentiate from classical codes $[n,k]$. Now suppose $C_{\_}1$ and $C_{\_}2$ are $[n,k_{\_}1]$ and $[n,k_{\_}2]$ classical linear codes with the property that $C_{\_}2\subseteq C_{\_}1$, and $C_{\_}1$ and $C^{\perp}_{\_}2$ both correct $q$ errors. The CSS code, constructed from $C_{\_}1$ and $C_{\_}2$, is then an $[[n,k_{\_}1-k_{\_}2]]$ quantum code that can correct arbitrary errors on up to $q$ qubits (though the code can actually correct up to $q$ bit-flip and $q$ phase-flip errors).
Codewords in the CSS code are of the form
|$\ket{w+C_{\_}2}\equiv\frac{1}{\sqrt{|C_{\_}2|}}\sum_{\_}{v\in C_{\_}2}\ket{w+v},$ (8)
where $w\in C_{\_}1$. Note that since $C_{\_}2\subseteq C_{\_}1$, $w+v\in C_{\_}1$. It is useful to observe the effect of a Hadamard transform on such a codeword,
$\frac{1}{\sqrt{|C^{\perp}_{\_}2|}}\sum_{\_}{z\in C^{\perp}_{\_}2}(-1)^{w\cdot z}\ket{z}.$
Bit-flip errors are described be an $n$-bit vector $e_{\_}1$ with 1s where bit-flips have occurred and 0s otherwise. Similarly, phase-flip errors are described by a vector $e_{\_}2$. Then the noisy state is
$\frac{1}{\sqrt{|C_{\_}2|}}\sum_{\_}{v\in C_{\_}2}(-1)^{(w+v)\cdot e_{\_}2}\ket{w+v+e_{\_}1},$
Error correction proceeds as follows. First, bit-flips are dealt with by applying the parity check matrix $H_{\_}{C_{\_}1}$ for the code $C_{\_}1$ and storing the result in ancilla qubits; measuring the ancilla qubits yields an error syndrome $H_{\_}{C_{\_}1}e_{\_}1$ that can be classically processed to determine an appropriate correction operation. After correction, the state is
$\frac{1}{\sqrt{|C_{\_}2|}}\sum_{\_}{v\in C_{\_}2}(-1)^{(w+v)\cdot e_{\_}2}\ket{w+v},$
For phase-flip errors, we apply a Hadamard transform to yield
$\frac{1}{\sqrt{|C^{\perp}_{\_}2|}}\sum_{\_}{z\in C^{\perp}_{\_}2}(-1)^{w\cdot z}\ket{z+e_{\_}2},$
and so we can now treat the phase-flip errors as bit-flips. We apply the parity check matrix $H_{\_}{C_{\_}2^{\perp}}$ for $C^{\perp}_{\_}2$ and obtain the error syndrome $H_{\_}{C_{\_}2^{\perp}}e_{\_}2$. After correction, another Hadamard transform is applied to return the state to the original codeword.
Let _$\mathcal{E}_{\_}X$ (_$\mathcal{E}_{\_}Z$) denote the set of error vectors that corresponds to bit-flip (phase-flip) errors affecting up to $q$ out of $n$ qubits:
_$\mathcal{E}_{\_}X=\{e\in\{0,1\}^{n}:\text{wt}(e)\leq q\}\quad\text{ and }\quad\mathcal{E}_{\_}Z=\{e^{\prime}\in\{0,1\}^{n}:\text{wt}(e^{\prime})\leq q\}.$ (9)
For a CSS code that corrects up to $q$ bit-flip and $q$ phase-flip errors (i.e., up to $q$ arbitrary errors), let _$\mathcal{E}_{\_}q$ denote the set of all correctable error vectors,
_$\mathcal{E}_{\_}q\coloneqq\{(e,e^{\prime}):e,e^{\prime}\in\{0,1\}^{n}\text{ and }\text{wt}(e)\leq q\text{ and }\text{wt}(e^{\prime})\leq q\}.$ (10)
We can easily see that
|_$|\mathcal{E}_{\_}q|=|\mathcal{E}_{\_}{X}|\cdot|\mathcal{E}_{\_}Z|=\left(\sum^{q}_{\_}{j=0}\binom{n}{j}\right)^{2}.$ (11)
With slight abuse of notation, we sometimes drop the $q$ subscript on _$\mathcal{E}_{\_}q$.
2.3.1 CSS code as a stabilizer code
Stabilizer codes [Got97] form an important class of quantum error correction codes. In this section, we use [Got97, NC10, Fuj15, Rof19] to briefly review the stabilizer formalism so that we may view CSS codes from this perspective.
The $n$-Pauli group $\mathcal{P}_{\_}n$ is defined as
$\mathcal{P}_{\_}n\coloneqq\{\pm 1,\pm i\}\times\{\mathds{1},X,Y,Z\}^{\otimes n}.$
An $n$-qubit stabilizer $\mathcal{S}$ is an abelian (i.e., commutative) subgroup of $\mathcal{P}_{\_}n$,
$\mathcal{S}\coloneqq\{S_{\_}i\in\mathcal{P}_{\_}n\,|-\mathds{1}\notin\mathcal{S}\text{ and }\forall S_{\_}i,S_{\_}j\in\mathcal{S},[S_{\_}i,S_{\_}j]=0\}.$
Elements of the maximum independent set $\mathcal{S}_{\_}g$ of the stabilizer are referred to as the stabilizer generators $g_{\_}i$, where independence is understood in the sense that no element of $\mathcal{S}_{\_}g$ can be expressed as a product of other elements in $\mathcal{S}_{\_}g$.
Given a stabilizer, the codespace $C_{\_}\mathcal{S}$ is the vector space that is stabilized by $\mathcal{S}$ in the sense that it consists of all $n$-qubit states which are fixed by every element of $\mathcal{S}$. Equivalently, $C_{\_}\mathcal{S}$ is the common +1 eigenspace of all elements of $\mathcal{S}$,
$C=\{\ket{\psi}:S_{\_}i\ket{\psi}=\ket{\psi},\,\,\forall S_{\_}i\in\mathcal{S}_{\_}g\}.$
A stabilizer is often described by its generators using the notation $\mathcal{S}_{\_}g=\langle g_{\_}1,\ldots,g_{\_}m\rangle$. Alternatively, the generators can be described by a check matrix. This is an $m\times 2n$ matrix where the $i$th row is constructed according to the following rules:
  • If $g_{\_}i$ contains an $I$ on the $j$th qubit, then the $j$th and $n+j$th column elements are 0.
  • If $g_{\_}i$ contains an $X$ on the $j$th qubit, then the $j$th column element is 1 and the $n+j$th column element is 0.
  • If $g_{\_}i$ contains a $Z$ on the $j$th qubit, then the $j$th column element is 0 and the $n+j$th column element is 1.
  • If $g_{\_}i$ contains a $Y$ on the $j$th qubit, then the $j$th and $n+j$th column elements are 1.
In the case of a CSS code, which is an example of a stabilizer code, the check matrix is of the form
$\begin{bmatrix}[c|c]H_{\_}{C^{\perp}_{\_}2}&0\\ 0&H_{\_}{C_{\_}1}\end{bmatrix}$ (12)
where $H$ is the parity check matrix for the corresponding classical linear code. So, from the parity check matrices $H_{\_}{C_{\_}1}$ and $H_{\_}{C_{\_}2^{\perp}}$, we can obtain the stabilizer generators. The $C_{\_}1$ code then contributes $n-k_{\_}1$ generators, and the $C^{\perp}_{\_}2$ code contributes $k_{\_}2$ generators, for a total of $n-k_{\_}1+k_{\_}2$ generators.
Error correction for a stabilizer code works in the following way. Let |$\ket{\psi}$ be our codeword and $E$ be some error. Then for each stabilizer generator $g_{\_}i$, $i=1,\ldots,m$, the state $E\ket{\psi}\ket{0}$ is mapped in the following way
[…]
After this, each of the $m$ ancilla qubits |$\ket{0}$ are measured and the results form the $m$-bit syndrome. A stabilizer generator will detect the error $E$ if it anti-commutes with $E$, in which case measurement of the ancilla qubit yields 1; if they commute, the measurement will yield 0. We can equivalently interpret this as: the generator $g_{\_}i$ detects the error $E$ if $E\ket{\psi}$ is in the $-1$ eigenspace of $g_{\_}i$, and the error will not be detected if $E\ket{g_{\_}i}$ is in the +1 eigenspace of $g_{\_}i$.
Given a stabilizer $\mathcal{S}$, we may have a set of operators that do not belong to $\mathcal{S}$ and yet commute with every element of $\mathcal{S}$. These operators, known as logical operators, can be interpreted as a means of computing on encoded states, as they act on codewords without moving them “out” of the codespace.
Note that every element of the stabilizer $\mathcal{S}$ has ±$\pm 1$ eigenvalues and recall that the codespace is defined as the common +1 eigenspace of the generators of $\mathcal{S}$. Now suppose there are $m$ stabilizer generators. Every stabilizer generator divides the Hilbert space of an $n$-qubit system into two orthogonal subspaces (the $+1$ eigenspace and the $-1$ eigenspace). Then, by commutativity of the stabilizer generators, the $m$ stabilizer generators divide the Hilbert space into $2^{m}$ orthogonal subspaces. Consequently, the codespace has dimension $2^{k}=2^{n-m}$. In the case of a CSS code, the codespace has dimension $2^{k}=2^{n-(n-k_{\_}1+k_{\_}2)}=2^{k_{\_}1-k_{\_}2}$.
3 Noise-tolerance for public-key quantum money
In this section, we formally define what we mean by a noise-tolerant public-key quantum money scheme and we fully describe our approaches to designing such a scheme. We start in section 3.1 with recalling how public-key quantum money was defined in [AC12]. Then, in section 3.2, we extend this definition to the noisy setting. Before describing our approaches to designing a noise-tolerant scheme (in section 3.4), we briefly outline alternative approaches in section 3.3, though these are not the main focus of the paper. The formal specification of our scheme is given in section 3.5.
3.1 Definitions
In this section, we recall how public-key quantum money is defined in [AC12]. Specifically, we focus on what Aaronson and Christiano call mini-schemes, since they show that a secure mini-scheme can be combined with a digital signature scheme to construct a full public-key quantum money scheme that is also secure. For further details on this construction, the reader is referred to [AC12].
Notably, a valid banknote in the context of 3.1 refers to a banknote as output by the algorithm Bank. For the noise-tolerant setting, we define a valid banknote differently (see 3.3).
Definition 3.1 (Mini-schemes [AC12]).
A (public-key) mini-scheme $\mathcal{M}$ consists of two polynomial-time quantum algorithms:
  • Bank, which takes as input a security parameter $0^{n}$, and probabilistically generates a banknote \$$\$=(s,\rho_{\_}s)$ where $s$ is a classical serial number, and $\rho_{\_}s$ is a quantum money state.
  • Ver, which takes as input an alleged banknote $\tilde{\$}$, and either accepts or rejects.
We say $\mathcal{M}$ has completeness error $\varepsilon$ if $\textsf{Ver}(\$)$ accepts with probability at least $1-\varepsilon$ for all valid banknotes \$. If $\varepsilon=0$, then $\mathcal{M}$ has perfect completeness. If, furthermore, $\rho_{\_}s=\ket{\psi_{\_}s}\bra{\psi_{\_}s}$ | is always a pure state, and Ver simply consists of a projective measurement onto the rank-1 subspace spanned by |$\ket{\psi_{\_}s}$, then we say that $\mathcal{M}$ is projective.
Let $\textsf{Ver}_{\_}2$ (the double verifier) take as input a single serial number $s$ as well as two (possibly entangled) states $\sigma_{\_}1$ and $\sigma_{\_}2$, and accept if and only if $\textsf{Ver}(s,\sigma_{\_}1)$ and $\textsf{Ver}(s,\sigma_{\_}2)$ both accept. We say that $\mathcal{M}$ has soundness error $\delta$ if, given any quantum circuit $\mathscr{C}$ of size $\textsf{poly}(n)$ (the counterfeiter), $\textsf{Ver}_{\_}2(s,\mathscr{C}(\$))$ accepts with probability at most $\delta$. Here, the probability is over the banknote \$ output by $\textsf{Bank}(0^{n})$, as well as the behaviour of $\textsf{Ver}_{\_}2$ and $\mathscr{C}$.
We call $\mathcal{M}$ secure if it has completeness error $\leq 1/3$ and negligible soundness error.
3.2 Defining a noise-tolerant mini-scheme
Put simply, the banknote in the money scheme of [AC12] is a subspace state |$\ket{A}$ and verification is performed by projecting onto the subspace spanned by |$\ket{A}$. If the state |$\ket{A}$ is corrupted by a Pauli error, then this state will be orthogonal to |$\ket{A}$ (see the end of the proof for 3.5) and hence rejected by this verification procedure. To extend the ideas of [AC12] to tolerate noise, we propose that verification of a banknote amounts to projecting onto the subspace spanned by the banknote and all its tolerated noisy variants (i.e. a subspace spanned by coset states, where the coset states are indexed by the set of tolerated error vectors). This idea, which is a natural extension of projective mini-schemes from 3.1, is formalized by the following definition.
Definition 3.2 (Noisy-projective mini-schemes).
Let $\mathcal{M}=(\textsf{Bank},\textsf{Ver})$ be a mini-scheme. We say that $\mathcal{M}$ is noisy-projective if Ver consists of a projective measurement onto the subspace spanned by $\{X^{e}Z^{e^{\prime}}\ket{\psi_{\_}s}\}_{\_}{(e,e^{\prime})\in\mathcal{E}_{\_}q}$, where _$\mathcal{E}_{\_}q$ is the set of tolerated error vectors for a scheme that can tolerate up to $q$ arbitrary errors on an $n$-qubit banknote (see equation 10).
Definition 3.3 (Valid banknotes for a noise-tolerant scheme).
Let $\mathcal{M}$ be a mini-scheme. In a noisy-projective $\mathcal{M}$, we say that a valid banknote is one that has been minted by the bank and then possibly affected by a tolerable error; that is, a pair consisting of a valid serial number $s$ and a state from the set $\{X^{e}Z^{e^{\prime}}\ket{\psi_{\_}s}\}_{\_}{(e,e^{\prime})\in\mathcal{E}_{\_}q}$.
Definition 3.4 (Security of a noisy-projective mini-scheme).
Let $\mathcal{M}$ be a mini-scheme. We say that a noisy-projective $\mathcal{M}$ has completeness error $\varepsilon$ if $\textsf{Ver}(\$)$ accepts with probability at least $1-\varepsilon$ for all valid banknotes \$currency-dollar\\$\$ (where valid is understood in the sense of 3.3). If $\varepsilon=0$, then $\mathcal{M}$ has perfect completeness.
Let $\textsf{Ver}_{\_}2$ (the double verifier) take as input a single serial number $s$ as well as two (possibly entangled) states $\sigma_{\_}1$ and $\sigma_{\_}2$, and accept if and only if $\textsf{Ver}(s,\sigma_{\_}1)$ and $\textsf{Ver}(s,\sigma_{\_}2)$ both accept. We say that $\mathcal{M}$ has soundness error $\delta$ if, given any quantum circuit $\mathscr{C}$ of size $\textsf{poly}(n)$ (the counterfeiter), $\textsf{Ver}_{\_}2(s,\mathscr{C}(\$))$ accepts with probability at most $\delta$. Here, the probability is over the banknote \$ output by $\textsf{Bank}(0^{n})$, as well as the behaviour of $\textsf{Ver}_{\_}2$ and $\mathscr{C}$.
When $q\geq 1$, we say that a noisy-projective $\mathcal{M}$ is $\epsilon$-noise-tolerant, where $\epsilon\coloneqq q/n$. We say that a noisy-projective $\mathcal{M}$ is $\delta$-secure if the scheme has soundness error $\delta$ and negligible completeness error.
3.3 Initial approaches to noise-tolerance
In this section, we briefly discuss two approaches to noise-tolerance that are at first intuitive but then prove to present obstacles. In section 3.3.1, we treat the banknote of the [AC12] scheme as already being a codeword of a CSS code. In section 3.3.2, we consider a different direction by adding a layer of error correction onto the [AC12] scheme.
3.3.1 Direct error correction
We can view the banknote subspace state in [AC12] as a codeword in a CSS code. Furthermore, from the perspective of a CSS code (with only one codeword), the verification procedure of [AC12] (checking membership in $A$ and $A^{\perp}$) can be thought of as error detection where one only checks for membership in the codespace (the common $+1$ eigenspace of the generators), as both procedures amount to projecting onto the subspace spanned by |$\ket{A}$. If we use this CSS code and the error correction procedure outlined in section 2.3 for noise-tolerance, then we must measure the generators of the code to determine whether the codeword has been affected by a correctable error vector. The immediate difficulty with this approach is that publishing the generators reveals the subspace used for constructing the banknote.
To formally see this, we first need to elaborate on why our CSS codes can only have a single codeword. If a CSS code has more than one codeword, then the verification procedure where one checks membership in the codespace allows all codewords to be accepted; in the noise-free quantum money setting, this is advantageous for a counterfeiter, as multiple states can now be accepted, as opposed to the single state |$\ket{A}$. In a noise-tolerant setting, this problem would only worsen.
To be precise in our characterization of applicable CSS codes, we are asking for a subspace $C$ such that $\dim(C)=n/2$ and that both $C$ and $C^{\perp}$ can correct up to $q$ errors. Then, setting $C_{\_}1=C$ and $C_{\_}2=C$ in the CSS code framework, we get a code that can correct up to $q$ arbitrary errors and where the codespace has $12^{k_{\_}1-k_{\_}2}=$ 1 element, as desired. The single codeword is then the subspace state
|$\ket{C}=\frac{1}{\sqrt{|C|}}\sum_{\_}{v\in C}\ket{v}.$
Returning to the task of noise-tolerance, we recall that measuring the generators of a code is how we determine whether a codeword has been affected by a correctable error vector. For an applicable CSS code, the check matrix (which describes the generators), is
$\begin{bmatrix}[c|c]H_{\_}{C^{\perp}}&0\\ 0&H_{\_}{C}\end{bmatrix}.$
Thus, if we were to publish a description of the generators for a banknote |$\ket{C}$, an adversary could determine the parity check matrices which in turn reveals the codespace $C$. It then seems that to use the generators for tolerating noise, one must “hide” a full description of them.
3.3.2 A layer of error correction
Now suppose that we take a more traditional approach by applying a layer of error correction on top of the oracle scheme of [AC12]. The process for this starts by choosing a stabilizer code and then encoding the banknote subspace state |$\ket{A}$. Given a state |$\ket{\psi}$, let |$\ket{\psi}_{\_}L$ denote the encoding of |$\ket{\psi}$. The natural way to prepare |$\ket{A}_{\_}L$ is to construct it from |$\ket{0}_{\_}L$, just as |$\ket{A}$ would be constructed from |$\ket{0}$.
The standard encoding procedure for a stabilizer code is to start by encoding the |$\ket{0}$ state via the following operation
|$\ket{0}_{\_}L\equiv\ket{0^{\otimes k}}_{\_}L=\frac{1}{N}\prod_{\_}{g_{\_}i\in\mathcal{S}_{\_}g}\left(\mathds{1}^{\otimes n}+g_{\_}i\right)\ket{0^{\otimes n}}$ (14)
where we recall that $\mathcal{S}_{\_}g$ denotes the set of stabilizer generators, and $N$ is for normalization. Note that with the codeword |$\ket{0}_{\_}L$, one can prepare the other codewords by applying logical operators to |$\ket{0}_{\_}L$. Now, given the circuit which transforms |$\ket{0}$ to |$\ket{A}$, one must now translate this circuit into one that operates on the encoded state |$\ket{0}_{\_}L$ to prepare |$\ket{A}_{\_}L$. Typically, the logical operators of a stabilizer code do not form a universal gate set, and so it is not immediately obvious how this can be done.
Assuming that |$\ket{A}_{\_}L$ can be prepared, one can take two directions. The first is to design a fault-tolerant version of the verification procedure which includes the queries to the classical oracle. The other approach is to perform error correction on |$\ket{A}_{\_}L$, decode it to |$\ket{A}$, and then apply the verification procedure. The latter approach is simpler but assumes that the verification procedure is noise-free. In either approach, a formal proof of security does not seem to immediately follow from the original security proof of [AC12]. The most obvious reason is that a codeword may be quite different from the state it encodes, meaning that the state held by an adversary, |$\ket{A}_{\_}L$, will be different from the state |$\ket{A}$ which the original security proof relies on.
3.4 Verification
In this section, we describe our noise-tolerant quantum money scheme. We present two approaches which we refer to as the the subset approach (section 3.4.1) and the subspace approach (section 3.4.2). In each section, we characterize the banknote and the verification procedure. As noted in section 1.2, we cover both approaches here because it is not immediately obvious which is preferable in the plain model (see also section 5), though the rest of the paper focuses primarily on the subset approach. The formal specification of our noisy-projective mini-scheme comes later in section 3.5. Then, in section 4, we analyze the security of our scheme.
3.4.1 The subset approach
We now describe the verification procedure that we use throughout the rest of the paper. To start, let $\mathcal{K}$ be the set of all subspaces $C\subseteq\mathbb{F}^{n}_{\_}2$ such that $\dim C=n/2$. Recalling equation 7, let $\mathcal{W}$ be the subset of $\mathcal{K}$ defined as
$\mathcal{W}\coloneqq\{C\in\mathcal{K}:d(C)\geq 2q+1\text{ and }d(C^{\perp})\geq 2q+1\},$ (15)
In other words, $\mathcal{W}$ consists of applicable CSS codes: $n/2$-dimensional subspaces $C$ of $\mathbb{F}^{n}_{\_}2$ such that, as classical linear codes, both $C$ and $C^{\perp}$ can each correct $q$ errors.
A freshly minted money state of our scheme will be a subspace state
|$\ket{C}=\frac{1}{\sqrt{|C|}}\sum_{\_}{v\in C}\ket{v},$ (16)
where $C\in\mathcal{W}$. This is similar to [AC12], where a freshly minted money state is also a subspace state except with $C\in\mathcal{K}$. The reason we must select our subspaces from $\mathcal{W}$ is to ensure that our verification operator, constructed below, is a projector; this is made clear in the proof of 3.5. We discuss the existence of such subspaces in the subspace approach (section 3.4.2), which also requires the $C\in\mathcal{W}$ condition.
When the state in equation 16 is corrupted by noise (a bit-flip error $X^{e}$ and a phase-flip error $Z^{e^{\prime}}$), it becomes a coset state
|$\ket{C_{\_}{e,e^{\prime}}}=X^{e}Z^{e^{\prime}}\ket{C}=\frac{1}{\sqrt{|C|}}\sum_{\_}{v\in C}(-1)^{v\cdot e^{\prime}}\ket{v+e},$
where $e,e^{\prime}$ are $n$-bit vectors. Suppose we wish to tolerate up to $q$ arbitrary errors; this means that we wish to tolerate the set _$\mathcal{E}_{\_}q$ defined by equation 10. Taking inspiration from CSS codes (section 2.3), we know that this tolerance can be achieved by tolerating up to $q$ bit-flip errors and up to $q$ phase-flip errors. Precisely, this means that our scheme should tolerate the sets _$\mathcal{E}_{\_}X,\mathcal{E}_{\_}Z$ defined by equation 9. To tolerate these two sets, we use classical oracles to check for membership in the subsets $C_{\_}{\mathcal{E}_{\_}X}$ and $C_{\_}{\mathcal{E}_{\_}Z}$ defined as
$C_{\_}{\mathcal{E}_{\_}X}\coloneqq\bigcup_{\_}{e\in\mathcal{E}_{\_}X}C+e\quad\text{ and }\quad C_{\_}{\mathcal{E}_{\_}Z}\coloneqq\bigcup_{\_}{e^{\prime}\in\mathcal{E}_{\_}Z}C^{\perp}+e^{\prime}$ (17)
Formally, tolerable bit-flip errors are checked by our first projector _$\mathbb{P}_{\_}{C}$ which is constructed as follows:
  1. 1.
    For the oracle $U_{\_}{C_{\_}{\mathcal{E}_{\_}X}}$, we introduce a control qubit $\ket{+}$ and apply the classical oracle controlled on $\ket{+}$.
  2. 2.
    We then measure the control qubit and accept if the outcome $\ket{-}$ is observed, and reject otherwise.
The projector _$\mathbb{P}_{\_}{C^{\perp}}$, which will check for tolerable phase-flip errors, is constructed in exactly the same except that the classical oracle $U_{\_}{C_{\_}{\mathcal{E}_{\_}Z}}$ is used and, for this projector to work, it will need to be preceded and followed by the Hadamard transform.
Verification is then performed by the following operator
$V=\mathsf{H}^{\otimes n}\mathbb{P}_{\_}{C^{\perp}}\mathsf{H}^{\otimes n}\mathbb{P}_{\_}{C}.$ (18)
The following lemma is a generalization of lemma 19 from [AC12]. It shows that our verification operator is a projector onto the subspace spanned by the money state |$\ket{C}$ and all its tolerated noisy variants.
Lemma 3.5.
The verification operator $V$, defined in equation 18, is the projector
$V=\sum_{\_}{(e,e^{\prime})\in\mathcal{E}}\ket{C_{\_}{e,e^{\prime}}}\bra{C_{\_}{e,e^{\prime}}}$ |, i.e., it projects onto the subspace $S$ spanned by {|$\{\ket{C_{\_}{e,e^{\prime}}}\}_{\_}{(e,e^{\prime})\in\mathcal{E}}$. In particular, the probability of acceptance is given by
$\Pr[V(\ket{\psi})=\text{accepts}]=\sum_{\_}{(e,e^{\prime})\in\mathcal{E}}|\braket{\psi}{C_{\_}{e,e^{\prime}}}|^{2}.$
Proof.
We first show that $V\ket{C_{\_}{e,e^{\prime}}}=\ket{C_{\_}{e,e^{\prime}}}$ for any $(e,e^{\prime})\in\mathcal{E}$.
$V\ket{C_{\_}{e,e^{\prime}}}$ $=V\frac{1}{\sqrt{|C|}}\sum_{\_}{v\in C}(-1)^{v\cdot e^{\prime}}\ket{v+e}$
$=\mathsf{H}^{\otimes n}\mathbb{P}_{\_}{C^{\perp}}\mathsf{H}^{\otimes n}\mathbb{P}_{\_}{C}\frac{1}{\sqrt{|C|}}\sum_{\_}{v\in C}(-1)^{v\cdot e^{\prime}}\ket{v+e}$
$=\mathsf{H}^{\otimes n}\mathbb{P}_{\_}{C^{\perp}}\mathsf{H}^{\otimes n}\frac{1}{\sqrt{|C|}}\sum_{\_}{v\in C}(-1)^{v\cdot e^{\prime}}\ket{v+e}$
$=\mathsf{H}^{\otimes n}\mathbb{P}_{\_}{C^{\perp}}\frac{1}{\sqrt{|C|}}\sum_{\_}{v\in C}(-1)^{v\cdot e^{\prime}}\frac{1}{\sqrt{2^{n}}}\sum_{\_}{z\in 2^{n}}(-1)^{(v+e)\cdot z}\ket{z}$
$=\mathsf{H}^{\otimes n}\mathbb{P}_{\_}{C^{\perp}}\frac{(-1)^{e\cdot e^{\prime}}}{\sqrt{2^{n}|C|}}\sum_{\_}{v\in C}\sum_{\_}{z\in 2^{n}}(-1)^{(v+e)\cdot(z+e^{\prime})}\ket{z}$
$=\mathsf{H}^{\otimes n}\mathbb{P}_{\_}{C^{\perp}}\frac{(-1)^{e\cdot e^{\prime}}}{\sqrt{2^{n}|C|}}\sum_{\_}{z^{\prime}\in 2^{n}}\sum_{\_}{v\in C}(-1)^{(v+e)\cdot z^{\prime}}\ket{z^{\prime}+e^{\prime}}$
$=\mathsf{H}^{\otimes n}\mathbb{P}_{\_}{C^{\perp}}\frac{(-1)^{e\cdot e^{\prime}}}{\sqrt{2^{n}/|C|}}\sum_{\_}{z^{\prime}\in C^{\perp}}(-1)^{z^{\prime}\cdot e}\ket{z^{\prime}+e^{\prime}}$
$=\mathsf{H}^{\otimes n}\frac{(-1)^{e\cdot e^{\prime}}}{\sqrt{|C^{\perp}|}}\sum_{\_}{z^{\prime}\in C^{\perp}}(-1)^{z^{\prime}\cdot e}\ket{z^{\prime}+e^{\prime}}$
$=\frac{1}{\sqrt{|C|}}\sum_{\_}{v\in C}(-1)^{v\cdot e^{\prime}}\ket{v+e}$
where we set $z^{\prime}\equiv z+e^{\prime}$ and used the result that if $z^{\prime}\in C^{\perp}$ then _$\sum_{\_}{v\in C}(-1)^{v\cdot z^{\prime}}=|C|$ and if $z^{\prime}\notin C^{\perp}$ then _$\sum_{\_}{v\in C}(-1)^{v\cdot z^{\prime}}=0$.
We now show that $V\ket{\psi}=0$ for all |$\ket{\psi}$ that are orthogonal to every tolerable noisy state |$\ket{C_{\_}{e,e^{\prime}}}$. We start by writing $\psi$ as
|$\ket{\psi}=\sum_{\_}{x\in 2^{n}}c_{\_}x\ket{x}.$
Then the orthogonality condition gives
00 0 $=\braket{\psi\,}{\,C_{\_}{e,e^{\prime}}}$
=(_$=\bigg{(}\sum_{\_}{x\in 2^{n}}c^{*}_{\_}x\bra{x},\frac{1}{\sqrt{|C|}}\sum_{\_}{v\in C}(-1)^{v\cdot e^{\prime}}\ket{v+e}\bigg{)}$
$\Longleftrightarrow 0$ =_$=\sum_{\_}{x\in C+e}c^{*}_{\_}x(-1)^{(x+e)\cdot e^{\prime}}$
$\Longleftrightarrow 0$ =_$=\sum_{\_}{v\in C}c_{\_}{v+e}(-1)^{v\cdot e^{\prime}}$
for all $(e,e^{\prime})\in\mathcal{E}$. Then,
$V\ket{\psi}$ $=\mathsf{H}^{\otimes n}\mathbb{P}_{\_}{C^{\perp}}\mathsf{H}^{\otimes n}\mathbb{P}_{\_}{C}\sum_{\_}{x\in 2^{n}}c_{\_}x\ket{x}$
$=\mathsf{H}^{\otimes n}\mathbb{P}_{\_}{C^{\perp}}\mathsf{H}^{\otimes n}\sum_{\_}{e\in\mathcal{E}_{\_}X}\bigg{(}\sum_{\_}{v\in C}c_{\_}{v+e}\ket{v+e}\bigg{)}$
$=\mathsf{H}^{\otimes n}\mathbb{P}_{\_}{C^{\perp}}\sum_{\_}{e\in\mathcal{E}_{\_}X}\bigg{(}\sum_{\_}{v\in C}c_{\_}{v+e}\bigg{(}\frac{1}{\sqrt{2^{n}}}\sum_{\_}{y\in 2^{n}}(-1)^{(v+e)\cdot y}\ket{y}\bigg{)}\bigg{)}$
$=\mathsf{H}^{\otimes n}\mathbb{P}_{\_}{C^{\perp}}\sum_{\_}{e\in\mathcal{E}_{\_}X}\bigg{(}\sum_{\_}{v\in C}c_{\_}{v+e}\bigg{(}\frac{1}{\sqrt{2^{n}}}\sum_{\_}{y^{\prime}\in 2^{n}}(-1)^{(v+e)\cdot(y^{\prime}+e^{\prime})}\ket{y^{\prime}+e^{\prime}}\bigg{)}\bigg{)}$
$=\mathsf{H}^{\otimes n}\mathbb{P}_{\_}{C^{\perp}}\sum_{\_}{e\in\mathcal{E}_{\_}X}\bigg{(}\sum_{\_}{v\in C}c_{\_}{v+e}\bigg{(}\frac{1}{\sqrt{2^{n}}}(-1)^{(v+e)\cdot e^{\prime}}\sum_{\_}{y^{\prime}\in 2^{n}}(-1)^{(v+e)\cdot y^{\prime}}\ket{y^{\prime}+e^{\prime}}\bigg{)}\bigg{)}$
$=\mathsf{H}^{\otimes n}\sum_{\_}{(e,e^{\prime})\in\mathcal{E}}\bigg{(}\sum_{\_}{v\in C}c_{\_}{v+e}\bigg{(}\frac{1}{\sqrt{2^{n}}}(-1)^{(v+e)\cdot e^{\prime}}\sum_{\_}{z\in C^{\perp}}\ket{z+e^{\prime}}\bigg{)}\bigg{)}$
$=\mathsf{H}^{\otimes n}\sum_{\_}{(e,e^{\prime})\in\mathcal{E}}\bigg{(}\bigg{(}\sum_{\_}{v\in C}c_{\_}{v+e}(-1)^{v\cdot e^{\prime}}\bigg{)}\bigg{(}\frac{(-1)^{e\cdot e^{\prime}}}{\sqrt{2^{n}}}\sum_{\_}{z\in C^{\perp}}\ket{z+e^{\prime}}\bigg{)}\bigg{)}$
$=0$
where $e^{\prime}\in\mathcal{E}_{\_}Z$.
Then $V$ is a sum of projectors, $V=\sum_{\_}{(e,e^{\prime})\in\mathcal{E}}\ket{C_{\_}{e,e^{\prime}}}\bra{C_{\_}{e,e^{\prime}}}$ |, which is itself a projector. Indeed, $V$ is a projector if and only if |$\ket{C_{\_}{e,e^{\prime}}}$ and |$\ket{C_{\_}{s,s^{\prime}}}$ are orthogonal whenever $(e,e^{\prime})\neq(s,s^{\prime})$.
$\braket{C_{\_}{e,e^{\prime}}}{C_{\_}{s,s^{\prime}}}=\bigg{(}\frac{1}{\sqrt{|C|}}\sum_{\_}{v\in C}(-1)^{v\cdot e^{\prime}}\bra{v+e},\frac{1}{\sqrt{|C|}}\sum_{\_}{z\in C}(-1)^{z\cdot s^{\prime}}\ket{z+s}\bigg{)}$
The above expression is indeed equal to zero for a fixed $C\in\mathcal{W}$ and when $(e,e^{\prime})\neq(s,s^{\prime})$. To see this, first observe that for a fixed $v$, $\braket{v+e\,}{\,z+s}=0$ unless $v+e=z+s$ for some $z\in C$, or rather, $z=v+e+s$, but this is impossible for $e\neq s$, since $v,z\in C$ and, because $d(C)\geq 2q+1$, we have $e,s,e+s\notin C$, and hence $z\in C$ while $v+e+s\notin C$. So, suppose that $e=s$ but, by the assumption $(e,e^{\prime})\neq(s,s^{\prime})$, we now have $e^{\prime}\neq s^{\prime}$. In the Fourier basis (apply the Hadamard transform) we get
$\bigg{(}\frac{(-1)^{e\cdot e^{\prime}}}{\sqrt{|C^{\perp}|}}\sum_{\_}{v\in C^{\perp}}(-1)^{v\cdot e}\bra{v+e^{\prime}},\frac{(-1)^{e\cdot e^{\prime}}}{\sqrt{|C^{\perp}|}}\sum_{\_}{z\in C^{\perp}}(-1)^{z\cdot s}\ket{z+s^{\prime}}\bigg{)}.$
Again, we get zero unless $z=v+e^{\prime}+s^{\prime}$ but $v,z\in C^{\perp}$ and, because $d(C^{\perp})\geq 2q+1$, we have $e^{\prime},s^{\prime},e^{\prime}+s^{\prime}\notin C^{\perp}$ and hence $z\in C^{\perp}$ while $v+e^{\prime}+s^{\prime}\notin C^{\perp}$. Thus, |$\ket{C_{\_}{e,e^{\prime}}}$ and |$\ket{C_{\_}{s,s^{\prime}}}$ are orthogonal because their error vectors are distinct and because $C\in\mathcal{W}$. ∎
As a consequence of the $C\in\mathcal{W}$ condition, we need the following lemma to prove theorem 4.7.
Lemma 3.6.
Let $f:\mathbb{F}^{n}_{\_}2\rightarrow\mathbb{F}^{n}_{\_}2$ be an invertible linear isometry. Then for any $C\in\mathcal{W}$, we have $f(C)\in\mathcal{W}$.
Proof.
A classical linear code with distance at least $12q+$ 1 is able to correct up to $q$ errors. Recall that the distance of a code $C$ is defined as
$d(C)\coloneqq\min_{\_}{x,y\in C,x\neq y}d(x,y)$
where $d(x,y)$ is the Hamming distance between $x$ and $y$. Since $f$ is an isometry (i.e., distance-preserving map) we have $d(f(x),f(y))=d(x,y)$ for all $x,y\in C$ and hence
$d(C)=d(f(C))$. Hence, $f(C)$ has the same distance and is able to correct the same number of errors. ∎
3.4.2 The subspace approach
In this approach to verification, let $C\in\mathcal{W}$, where $\mathcal{W}$ is defined by equation 15 (i.e., $C$ is an applicable CSS code, as described in section 3.3.1). Then observe that an error syndrome of all zeros indicates a codeword that has been unaffected by noise (see remark 3.7); put another way, the codeword belongs to the common $+1$ eigenspace of all the generators. If the error syndrome contains, for instance, two $1$’s, then we can interpret this as the banknote state belonging to the intersection of all $+1$ eigenspaces and two $-1$ eigenspaces. From this perspective, each valid syndrome corresponds to a particular intersection of eigenspaces.
Remark 3.7.
In general, an error syndrome of all zeros indicates that the codeword has been unaffected by noise or affected by undetectable noise. To elaborate, let $L$ be a logical operator of a stabilizer code and let |$\ket{\psi}$ be a codeword. Recall that because $L$ commutes with all elements of the stabilizer, $L\ket{\psi}$ will also produce a syndrome of all zeros (i.e., $L\ket{\psi}$ is still in the codespace). So, it is not necessarily true that a syndrome of all zeros correlates to a noiseless codeword. However, in the case of an applicable CSS code where there is only a single codeword (i.e., no logical operators), we can say that the all-zero syndrome implies a noiseless codeword.
Formally, let $\mathfrak{S}$ denote the set of good syndromes (i.e., if $s\in\mathfrak{S}$, then the error that produced $s$ is correctable). Under the stabilizer formalism, every $s\in\mathfrak{S}$ corresponds to a particular intersection of eigenspaces of the generators. Denote such a subspace as $Q_{\_}s$, where $s\in\mathfrak{S}$ indexes the subspace. Then for each $s\in\mathfrak{S}$, we want to check membership in $Q_{\_}s$.
In the context of a CSS code, the syndrome is a concatenation $s=s^{X}\|s^{Z}$, where $s^{X}$ ($s^{Z}$) denotes the syndrome that arises from bit-flip (phase-flip) errors. So, for a given $s\in\mathfrak{S}$, instead of checking membership in the single subspace $Q_{\_}s$, one checks for membership in two subspaces, $Q_{\_}{s^{X}}$ and $Q_{\_}{s^{Z}}$. Let $\mathfrak{S}^{X},\mathfrak{S}^{Z}$ denote, respectively, the set of syndromes that correspond to correctable bit-flip and phase-flip errors. Then let $Q_{\_}{\mathfrak{S}^{X}},Q_{\_}{\mathfrak{S}^{Z}}$ denote the following unions of subspaces,
$Q_{\_}{\mathfrak{S}^{X}}\coloneqq\bigcup_{\_}{s^{X}\in\mathfrak{S}^{X}}Q_{\_}{s^{X}}\quad\text{ and }\quad Q_{\_}{\mathfrak{S}^{Z}}\coloneqq\bigcup_{\_}{s^{Z}\in\mathfrak{S}^{Z}}Q^{\mathsf{H}}_{\_}{s^{Z}}.$
where $Q^{\mathsf{H}}_{\_}{s^{Z}}$ denotes $Q_{\_}{s^{Z}}$ in the Hadamard basis. The first projector _$\mathbb{P}_{\_}{C}$ is then constructed as follows:
  1. 1.
    For the oracle $U_{\_}{Q_{\_}{\mathfrak{S}^{X}}}$, we introduce a control qubit $\ket{+}$ and apply the classical oracle controlled on $\ket{+}$.
  2. 2.
    We then measure the control qubit and accept if the outcome $\ket{-}$ is observed, and reject otherwise.
The process for phase-flip errors is only slightly different: repeat the above steps using $Q_{\_}{\mathfrak{S}^{Z}}$ to construct the second projector _$\mathbb{P}_{\_}{C^{\perp}}$. The final verification operator would then be given by $V=\mathsf{H}^{\otimes n}\mathbb{P}_{\_}{C^{\perp}}\mathsf{H}^{\otimes n}\mathbb{P}_{\_}{C}$.
With this $V$, 3.5 also holds. It is worth noting that if we do not use an applicable CSS code, then 3.5 does not necessarily hold, since we can no longer guarantee that $V$ only accepts a codeword and its tolerated noisy variants. A codeword affected by a logical operator (i.e., a highly corrupted codeword) could be accepted, but this possibility is eliminated by using an applicable CSS code.
To end this section, we remark that the existence of applicable CSS codes, $C\in\mathcal{W}$, is guaranteed by the Gilbert-Varshamov bound for CSS codes. This bound says that an $[[n,k]]$ quantum code which corrects up to $q$ arbitrary errors exists for some $k$ such that
$\frac{k}{n}\geq 1-2H\left(\frac{2q}{n}\right)$ (19)
where $H(x)=-x\log(x)-(1-x)\log(1-x)$ is the binary Shannon entropy. In our case, we have $k=0$.
Figure 1: A plot of the right-hand-side of equation 19 for various choices of $q$. For a fixed choice of $q$, equation 19 guarantees the existence of an applicable CSS codes as long as $n$ is chosen such that the corresponding point on the plot is below the horizontal axis.
As an example of an applicable CSS code, consider the following $[[6,3]]$ classical linear code $C$,
$G_{\_}C=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\\ 0&1&1\\ 1&1&0\\ 1&0&1\end{pmatrix},\quad H_{\_}C=\begin{pmatrix}0&1&1&1&0&0\\ 1&1&0&0&1&0\\ 1&0&1&0&0&1\end{pmatrix}.$
Since the distance of this code is $d=3$ and the inequality $d\geq 2q+1$ is satisfied for $q=1$, this code can correct a single error. The dual code $C^{\perp}$ has the following generator and parity check matrix,
$G_{\_}{C^{\perp}}=H_{\_}C^{T}=\begin{pmatrix}0&1&1\\ 1&1&0\\ 1&0&1\\ 1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix},\quad H_{\_}{C^{\perp}}=G_{\_}C^{T}=\begin{pmatrix}1&0&0&0&1&1\\ 0&1&0&1&1&0\\ 0&0&1&1&0&1\end{pmatrix}.$
The $C^{\perp}$ code also has a distance of $d=3$ and can thus also correct a single error. Hence, as a CSS code with $C_{\_}1,C_{\_}2=C$ we can correct a single arbitrary error. The single codeword of this CSS code is given by
[…]
3.5 Formal specification of the scheme
We must now clarify how the bank actually provides access to the oracles $U_{\_}{C_{\_}{\mathcal{E}_{\_}X}}$ and $U_{\_}{C_{\_}{\mathcal{E}_{\_}Z}}$ in the subset approach. In the final noisy-projective mini-scheme $\mathcal{M}=(\textsf{Bank}_{\_}\mathcal{M},\textsf{Ver}_{\_}\mathcal{M})$, the bank, verifier, and counterfeiter will all have access to a single classical oracle $U$, which consists of the four components described below. In section 3.5.1, we show how this specification can be amended for the purpose of using conjugate coding states.
  • A banknote generator $\mathcal{G}(r)$ which takes as input a random string $r\in\{0,1\}^{n}$, and outputs a set of linearly independent generators $\langle C_{\_}r\rangle=\{x_{\_}1,\ldots,x_{\_}{n/2}\}$ for a subspace $C_{\_}r\in\mathcal{W}$, as well as a unique $3n$-bit serial number $z_{\_}r\in\{0,1\}^{3n}$. The function $\mathcal{G}$ is chosen uniformly at random, subject to the constraint that the serial numbers are all distinct.
  • A serial number checker $\mathcal{Z}(z)$, which outputs 1 if $z=z_{\_}r$ is a valid serial number for some $\langle C_{\_}r\rangle$, and 0 otherwise.
  • A primal subset tester $\mathcal{T}_{\_}{\text{primal}}$ which takes an input of the form |$\ket{s}\ket{x}$, applies $U_{\_}{C_{\_}{\mathcal{E}_{\_}X}}$ to |$\ket{x}$ if $z=z_{\_}r$ is a valid serial number for some $\langle C_{\_}r\rangle$, and does nothing otherwise.
  • A dual subset tester $\mathcal{T}_{\_}{\text{dual}}$, identical to $\mathcal{T}_{\_}{\text{primal}}$ except that it applies $U_{\_}{C_{\_}{\mathcal{E}_{\_}Z}}$ instead of $U_{\_}{C_{\_}{\mathcal{E}_{\_}X}}$.
Then $\mathcal{M}=(\textsf{Bank}_{\_}\mathcal{M},\textsf{Ver}_{\_}\mathcal{M})$ is defined as follows:
[…]
  • $\textsf{Ver}_{\_}\mathcal{M}(\tilde{\$})$ first uses $\mathcal{Z}$ to check that $\tilde{\$}$ has the form $(z,\rho)$, where $z=z_{\_}r$ is a valid serial number. If so, then it uses $\mathcal{T}_{\_}{\text{primal}}$ and $\mathcal{T}_{\_}{\text{dual}}$ to apply $V_{\_}{C_{\_}r}=\mathsf{H}^{\otimes n}\mathbb{P}_{\_}{C_{\_}r^{\perp}}\mathsf{H}^{\otimes n}\mathbb{P}_{\_}{C_{\_}r}$, and accepts if and only if $V_{\_}{C_{\_}r}(\rho)$ accepts.
Remark 3.8.
The formal specification for the subspace approach is similar except that $U_{\_}{Q_{\_}{\mathfrak{S}^{X}}}$ and $U_{\_}{Q_{\_}{\mathfrak{S}^{Z}}}$ are used, respectively, in place of $U_{\_}{C_{\_}{\mathcal{E}_{\_}X}}$ and $U_{\_}{C_{\_}{\mathcal{E}_{\_}Z}}$.
3.5.1 Using conjugate coding states
Given that conjugate coding states are practically simple to prepare, it would be ideal if the money state of our scheme could be prepared with conjugate coding states, rather than directly preparing a subspace state |$\ket{A}$. In this section, we show how this can be done with 2.1 and, consequently, alter the banknote generator $\mathcal{G}(r)$ and $\textsf{Bank}_{\_}{\mathcal{M}}$.
Recalling 2.1, we see that generating a uniformly random coset state |$\ket{A_{\_}{t,t^{\prime}}}$ over subspaces $A\subseteq\mathbb{F}^{n}_{\_}2$ with dimension $n/2$ is equivalent to generating a basis =$\mathcal{B}=\{u_{\_}1,\ldots,u_{\_}n\}$ and $x,\theta\in\{0,1\}^{n}$, such that $\theta$ has Hamming weight $n/2$, uniformly at random. This leads to the following procedure:
  1. 1.
    Choose $n\geq 2$ to be even. Select $x\in\{0,1\}^{n}$ uniformly at random and select $\theta\in\{0,1\}^{n}$, such that $\theta$ has Hamming weight $n/2$, uniformly at random. Then form the state |$\ket{x}_{\_}\theta$.
  2. 2.
    Select a uniformly random basis =$\mathcal{B}=\{u_{\_}1,\ldots,u_{\_}n\}$ of $\mathbb{F}^{n}_{\_}2$.
  3. 3.
    Set $T=\{i:\theta_{\_}i=0\}$ and then set $A=\text{Span}\{u_{\_}i:\theta_{\_}i=1\}=\text{Span}\{u_{\_}i:i\in\overline{T}\}$. Then $A$ is a uniformly random subspace of $\mathbb{F}^{n}_{\_}2$ of dimension $n/2$.
  4. 4.
    Set $t=\sum_{\_}{i\in T}x_{\_}iu_{\_}i$ and $t^{\prime}=\sum_{\_}{i\in\overline{T}}x_{\_}iu^{i}$.
  5. 5.
    By 2.1, applying $U_{\_}{\mathcal{B}}$ to |$\ket{x}_{\_}\theta$ yields |$\ket{A_{\_}{t,t^{\prime}}}$.
In our scheme, a freshly minted money state (not yet corrupted by noise) is a subspace state |$\ket{C}$ with $C\in\mathcal{W}$. To generate a uniformly random subspace state |$\ket{C}$, we need $t=t^{\prime}=0$ in the above construction. A non-trivial choice of $x$ cannot achieve this without contradicting the fact that $\{u_{\_}1,\ldots,u_{\_}n\}$ forms a basis, thus we must have $x=0$. So, to generate a uniformly random subspace state under the above construction, we must generate |$\ket{0}_{\_}\theta$.
To then incorporate conjugate coding states into the formal specification of our scheme, we must alter the behaviour of the banknote generator $\mathcal{G}(r)$ and $\textsf{Bank}_{\_}\mathcal{M}$.
  • The banknote generator $\mathcal{G}(r)$ takes as input a random string $r\in\{0,1\}^{n}$, and outputs a basis _$\mathcal{B}_{\_}r=\{u_{\_}1,\ldots,u_{\_}n\}$ and a $\theta_{\_}r\in\{0,1\}^{n}$ with Hamming weight $n/2$ such that the subspace specified by $\theta_{\_}r$ belongs to $\mathcal{W}$, as well as a unique $3n$-bit serial number $z_{\_}r\in\{0,1\}^{3n}$. The function $\mathcal{G}$ is chosen uniformly at random, subject to the constraint that the serial numbers are all distinct.
[…]
Intuitively, using conjugate coding states in the minting of a banknote does not affect the security of the scheme since a counterfeiter in both variants of the scheme only receives a subspace state and access to the oracles. Formally, this change is handled in theorem 4.9.
4 Security
To prove security, we follow the same approach of [AC12], changing the proofs as necessary. To start, we look at one of the crucial components from [AC12] that was used for proving security: the amplification of counterfeiters theorem. Intuitively, the theorem shows that a counterfeiter that can copy a banknote with small success can be used to copy a banknote almost perfectly; hence, it suffices to prove security against counterfeiters that can copy almost perfectly. The analogous statement for noisy-projective mini-schemes is given as theorem 4.1 below. The proof of this theorem in [AC12] essentially starts with the following equality
$\Pr[V(\rho)=\text{accepts}]=F(\rho,S)^{2},$ (20)
where $F(\rho,S)$ is the fidelity of $\rho$ with the subspace $S$ that the verification algorithm $V$ projects to. Unlike [AC12], our space $S$ is not 1-dimensional (see 3.5), though we can still rely on the proof in [AC12] to get the same complexity bound. However, we will later see how having a “larger” $S$ will lead to a trade-off between security and noise-tolerance.
Theorem 4.1 (Amplification of counterfeiters).
[…]
$F(\mathscr{C}(\$),S^{\otimes 2})\geq\sqrt{\varepsilon}$
and hence
$\Pr[V^{\otimes 2}(\mathscr{C}(\$))=\text{accepts}]\geq\varepsilon.$
Then for all $\delta>0$, there is also a modified counterfeiter $\mathscr{C}^{\prime}$ (depending only on $\varepsilon$ and $\delta$, not \$currency-dollar\\$\$), which makes
$\mathcal{O}\left(\frac{\log 1/\delta}{\sqrt{\varepsilon}(\sqrt{\varepsilon}+\delta^{2})}\right)$ (21)
queries to $\mathscr{C},\mathscr{C}^{-1}$, and $V$ and which satisfies
$F(\mathscr{C}^{\prime}(\$),S^{\otimes 2})\geq 1-\delta$
which implies
$\Pr[V^{\otimes 2}(\mathscr{C}^{\prime}(\$))=\text{accepts}]\geq(1-\delta)^{2}.$
Proof.
We first note that
$\Pr[V(\rho)=\text{accepts}]$ $=\operatorname{Tr}(V\rho)$
$=\operatorname{Tr}\left(\sum_{\_}{(e,e^{\prime})\in\mathcal{E}}\ket{C_{\_}{e,e^{\prime}}}\bra{C_{\_}{e,e^{\prime}}}\rho\right)$
=_$=\sum_{\_}{(e,e^{\prime})\in\mathcal{E}}\operatorname{Tr}\left(\ket{C_{\_}{e,e^{\prime}}}\bra{C_{\_}{e,e^{\prime}}}\rho\right)$
=_$=\sum_{\_}{(e,e^{\prime})\in\mathcal{E}}\bra{C_{\_}{e,e^{\prime}}}\rho\ket{C_{\_}{e,e^{\prime}}}$
$=F(\rho,S)^{2}.$
Then, for the case of verifying two banknotes supplied by the counterfeiter $\mathscr{C}$, we have,
$\Pr[V^{\otimes 2}(\mathscr{C}(\rho))=\text{accepts}]$ =_$=\sum_{\_}{(e,e^{\prime}),(t,t^{\prime})\in\mathcal{E}}\bra{C_{\_}{e,e^{\prime}}}\otimes\bra{C_{\_}{t,t^{\prime}}}\mathscr{C}(\rho)\ket{C_{\_}{e,e^{\prime}}}\otimes\ket{C_{\_}{t,t^{\prime}}}$
=_$=\sum_{\_}{(e,e^{\prime}),(t,t^{\prime})\in\mathcal{E}}F(\mathscr{C}(\rho),\ket{C_{\_}{e,e^{\prime}}}\otimes\ket{C_{\_}{t,t^{\prime}}})^{2}$
$=F(\mathscr{C}(\rho),S^{\otimes 2})^{2}$
By assumption of the theorem,
$F(\mathscr{C}(\rho),S^{\otimes 2})\geq\sqrt{\varepsilon}$
and thus we indeed have
$\Pr[V^{\otimes 2}(\mathscr{C}(\rho))=\text{accepts}]\geq\varepsilon.$
Put simply, one can now apply a fixed-point Grover search with $\mathscr{C}(\rho)$ as the initial state and $S^{\otimes 2}$ as the target subspace to produce a state $\rho^{\prime}$ such that $F(\rho^{\prime},S^{\otimes 2})\geq 1-\delta$. As shown in [AC12] (see Section 2.3 and the proof of Theorem 13 for details), this can be done in
$\mathcal{O}\left(\frac{\log 1/\delta}{\sqrt{\varepsilon}(\sqrt{\varepsilon}+\delta^{2})}\right)$ (22)
Grover iterations. Each iteration can be implemented using $\mathcal{O}(1)$ queries to $\mathscr{C},\mathscr{C}^{-1}$, and $V$. Thus, the total number of queries to $\mathscr{C},\mathscr{C}^{-1}$, and $V$ is also given by equation 22. Then the probability of $\rho^{\prime}$ being accepted is
$\Pr[V^{\otimes 2}(\rho^{\prime})=\text{accepts}]$ =_$=\sum_{\_}{(e,e^{\prime}),(t,t^{\prime})\in\mathcal{E}}F(\rho^{\prime},\ket{C_{\_}{e,e^{\prime}}}\otimes\ket{C_{\_}{t,t^{\prime}}})^{2}$
$=F(\rho^{\prime},S^{\otimes 2})^{2}$
$\geq(1-\delta)^{2}.$
4.1 The inner-product adversary method
Put simply, proving security amounts to proving that a counterfeiter, who possesses a copy of |$\ket{\psi}$, needs to make exponentially many queries to the verification algorithm to prepare |$\ket{\psi_{\_}e}\ket{\psi_{\_}t}$, where |$\ket{\psi_{\_}e}$ and |$\ket{\psi_{\_}t}$ are possibly different noisy copies of |$\ket{\psi}$. For simplicity, let us forget about noise-tolerance so that the counterfeiter is now trying to prepare |$\ket{\psi}\ket{\psi}$. Suppose that |$\ket{\psi}$ and |$\ket{\phi}$ are two possible quantum money states. Consider a counterfeiting algorithm $\mathscr{C}$ that is executed in parallel to clone |$\ket{\psi}$ and |$\ket{\phi}$. If, say, $\braket{\psi}{\phi}=1/2$, and if the counterfeiting algorithm $\mathscr{C}$ succeeds perfectly (i.e., it maps |$\ket{\psi}$ to |$\ket{\psi}^{\otimes 2}$ and |$\ket{\phi}$ to |$\ket{\phi}^{\otimes 2}$), then $\bra{\psi}^{\otimes 2}\ket{\phi}^{\otimes 2}=1/4$, meaning that $\mathscr{C}$ must decrease the inner product by at least 1/4. But the analysis of [AC12] shows that the average inner product can decrease by at most $1/\textsf{exp}(n)$ from a single query, which results in $2^{\Omega(n)}$ queries being needed.
This analysis is done using the inner-product adversary method developed in [AC12] (which is an adaptation of the quantum adversary method in [Amb02]). Let us now recall the basic ideas of the method (for further details, see [AC12]). The idea is to get an upper bound on the progress a quantum algorithm $\mathscr{C}$ can make at distinguishing pairs of oracles, as a result of a single query. To make this a bit more formal, let |$\ket{\Psi^{U}_{\_}t}$ be $\mathscr{C}$’s state after $t$ queries. For now, assume that |$\ket{\Psi^{U}_{\_}0}=\ket{\Psi^{V}_{\_}0}$ for all oracles $U$ and $V$. After the final query $T$, we must have, say, $|\braket{\Psi^{U}_{\_}T}{\Psi^{V}_{\_}T}|\leq 1/2$. Thus, if we are able to show that $|\braket{\Psi^{U}_{\_}T}{\Psi^{V}_{\_}T}|$ | can decrease by at most $\varepsilon$ as the result of a single query, then it follows that $\mathscr{C}$ must make $\Omega(1/\varepsilon)$ queries.
The difficulty, overcome in [AC12], is that in the quantum money framework, $\mathscr{C}$ starts with a valid banknote and hence it is not generally true that |$\ket{\Psi^{U}_{\_}0}=\ket{\Psi^{V}_{\_}0}$. This starting point is advantageous for $\mathscr{C}$ since it can allow them to do much better in decreasing $|\braket{\Psi^{U}_{\_}T}{\Psi^{V}_{\_}T}|$ | as a result of a single query. The solution in [AC12] is to choose a distribution $\mathcal{D}$ over oracle pairs $(U,V)$ and then analyze the expected inner product
$\mathop{\mathbb{E}}_{\_}{(U,V)\sim\mathcal{D}}\left[|\braket{\Psi^{U}_{\_}t}{\Psi^{V}_{\_}t}|\right]$
and how much it can decrease as the result of a single query to $U$ or $V$. Analysis shows that, although $\mathscr{C}$ can succeed in significantly reducing $|\braket{\Psi^{U}_{\_}T}{\Psi^{V}_{\_}T}|$ | for some oracle pairs $(U,V)$, they cannot do so for most pairs.
4.1.1 Formal description
This section, which gives the formal description of the inner-product adversary method, is taken almost directly from [AC12]. Let $\mathcal{O}$ denote a set of quantum oracles acting on $n$ qubits each. For each $U\in\mathcal{O}$, assume there exists a subspace $S_{\_}U\subseteq\mathbb{C}^{2n}$ such that
  1. (i)
    $U\ket{\psi}=-\ket{\psi}$ for all |$\ket{\psi}\in S_{\_}U$
  2. (ii)
    $U\ket{\eta}$ for all |$\ket{\eta}\in S^{\perp}_{\_}U$.
Let $R\subset\mathcal{O}\times\mathcal{O}$ be a symmetric binary relation on $\mathcal{O}$, with the properties that
  1. (i)
    $(U,U)\notin R$ for all $U\in\mathcal{O}$, and
  2. (ii)
    for every $U\in\mathcal{O}$ there exists a $V\in\mathcal{O}$ such that $(U,V)\in R$.
Suppose that for all $U\in\mathcal{O}$ and all |$\ket{\eta}\in S^{\perp}_{\_}U$, we have
$\mathop{\mathbb{E}}_{\_}{V:(U,V)\in R}\Big{[}F(\ket{\eta},S_{\_}V)^{2}\Big{]}\leq\varepsilon,$
where $F(\ket{\eta},S_{\_}V)=\max_{\_}{\ket{\psi}\in S_{\_}V}|\braket{\eta}{\psi}|$ | is the fidelity between |$\ket{\eta}$ and $S_{\_}V$. Let $\mathscr{C}$ be a quantum oracle algorithm, and let $\mathscr{C}^{U}$ denote $\mathscr{C}$ run with the oracle $U\in\mathcal{O}$. Suppose $\mathscr{C}^{U}$ begins in the state |$\ket{\Psi^{U}_{\_}0}$ (possibly dependent on $U$). Let |$\ket{\Psi^{U}_{\_}t}$ denote the state of $\mathscr{C}^{U}$ immediately after the $t^{th}$ query. Also, define a progress measure $p_{\_}t$ by
$p_{\_}t\coloneqq\mathop{\mathbb{E}}_{\_}{V:(U,V)\in R}\Big{[}|\braket{\Psi^{U}_{\_}t}{\Psi^{V}_{\_}t}|\Big{]}.$
The following lemma bounds how much $p_{\_}t$ can decrease as the result of a single query. See [AC12] for a proof.
Lemma 4.2 (Bound on progress rate [AC12]).
$p_{\_}t\geq p_{\_}{t-1}-4\sqrt{\varepsilon}.$
As a result of 4.2, we have
Lemma 4.3 (Inner-product adversary method [AC12]).
Suppose that initially $|\braket{\Psi^{U}_{\_}0}{\Psi^{V}_{\_}0}|\geq~{}c$ for $(U,V)\in R$, whereas by the end we need $|\braket{\Psi^{U}_{\_}T}{\Psi^{V}_{\_}T}|\leq d$ for all $(U,V)\in R$. Then $\mathscr{C}$ must make
$T=\Omega\bigg{(}\frac{c-d}{\sqrt{\varepsilon}}\bigg{)}$
oracle queries.
4.1.2 Applying the method
For much of the security proof, we suppose that the counterfeiter only has access to the oracles $(U_{\_}{C_{\_}{\mathcal{E}_{\_}X}},U_{\_}{C_{\_}{\mathcal{E}_{\_}Z}})$ as opposed to the oracles in the noisy-projective mini-scheme specified in section 3.5. The generalization to the full scheme will be done later in theorem 4.9.
Our starting point for the analysis will be to think of the pair of oracles $(U_{\_}{C_{\_}{\mathcal{E}_{\_}X}},U_{\_}{C_{\_}{\mathcal{E}_{\_}Z}})$ as a single oracle. In the case of [AC12], they have the pair of oracles $(U_{\_}A,U_{\_}{A^{\perp}})$ and the corresponding single oracle is constructed in the following way. First, they consider a subset $A^{*}\subset\{0,1\}^{n+1}$ defined by
$A^{*}\coloneqq(0,A)\cup(1,A^{\perp}).$
Then they let $S_{\_}{A^{*}}$ denote the subspace of $\mathbb{C}^{2n+1}$ that is spanned by basis states |$\ket{x}$ such that $x\in A^{*}$. This allows their collection of oracles $(U_{\_}A,U_{\_}{A^{\perp}})$ to be thought of as a single oracle $U_{\_}{A^{*}}$ which satisfies
[…]
To generalize this to our setting, first suppose that our scheme only tolerates the errors =$\mathcal{E}=\{(e,e^{\prime}),(t,t^{\prime})\}$. Then the subset $C^{*}\subset\{0,1\}^{n+2}$, defined by
$C^{*}\coloneqq(00,C+e)\cup(01,C^{\perp}+e^{\prime})\cup(10,C+t)\cup(11,C^{\perp}+t^{\prime}),$ (23)
would be a generalization of $A^{*}$. Let us also introduce the following notation which naturally generalizes for a larger set of error vectors,
$C^{*}_{\_}{\mathcal{E}_{\_}X}$ $\coloneqq(00,C+e)\cup(10,C+t)$
$C^{*}_{\_}{\mathcal{E}_{\_}Z}$ $\coloneqq(01,C^{\perp}+e^{\prime})\cup(11,C^{\perp}+t^{\prime}).$
In general, every error vector from _$\mathcal{E}_{\_}X$ and _$\mathcal{E}_{\_}Z$ corresponds to a term in the union that defines $C^{*}$. So, we would have $2|\mathcal{E}_{\_}X|$ (i.e., |_$|\mathcal{E}_{\_}X|+|\mathcal{E}_{\_}Z|$) terms in the union, meaning $C^{*}\subset\{0,1\}^{n+k}$ where $k$ is such that
$2^{k}=2|\mathcal{E}_{\_}X|\Leftrightarrow k=1+\log(|\mathcal{E}_{\_}X|)$ (24)
Now let $S_{\_}{C^{*}}$ be the subspace of $\mathbb{C}^{2^{n+k}}$ that is spanned by basis states |$\ket{x}$ such that $x\in C^{*}$. Then our pair of oracles $(U_{\_}{C_{\_}{\mathcal{E}_{\_}X}},U_{\_}{C_{\_}{\mathcal{E}_{\_}Z}})$ corresponds to the single oracle $U_{\_}{C^{*}}$ which satisfies
[…]
A notable consequence of this generalization to the noisy setting is that a single query to $U_{\_}{C_{\_}{\mathcal{E}_{\_}X}}$ equates to |_$|\mathcal{E}_{\_}X|$ queries to $U_{\_}{C^{*}}$, and likewise for $U_{\_}{C_{\_}{\mathcal{E}_{\_}Z}}$. To see this, observe the example of $C^{*}$ in equation 23 where a query to $U_{\_}{C_{\_}{\mathcal{E}_{\_}X}}$ is achieved by preparing ancilla qubits set to |$\ket{00}$ and |$\ket{10}$ and then querying $U_{\_}{C^{*}}$ with each of these ancilla qubits. For security, this means that the lower bound for the number of queries to $(U_{\_}{C_{\_}{\mathcal{E}_{\_}X}},U_{\_}{C_{\_}{\mathcal{E}_{\_}Z}})$ is the lower bound for queries to $U_{\_}{C^{*}}$ divided by $2|\mathcal{E}_{\_}X|$. We return to this point in corollary 4.8.
The following theorem shows that, for the case of perfect counterfeiting, exponentially many queries are needed. In [AC12], perfect counterfeiting meant that the counterfeiter, who possessed a copy of |$\ket{C}$, had to prepare |$\ket{C}\otimes\ket{C}$. In our noise-tolerant setting, the counterfeiter has more flexibility by only needing to prepare noisy copies of |$\ket{C}$. This difference adds some extra steps to the proof but the lower bound on the needed queries remains the same as the analogous result in [AC12].
Theorem 4.4 (Lower bound for perfect counterfeiting).
Given one copy of |$\ket{C}$, as well as oracle access to $U_{\_}{C^{*}}$, a counterfeiter needs $\Omega\left(2^{n/4}\right)$ queries to prepare |$\ket{C_{\_}{e,e^{\prime}}}\otimes\ket{C_{\_}{t,t^{\prime}}}$, where $(e,e^{\prime}),(t,t^{\prime})\in\mathcal{E}$, with certainty (for a worst-case |$\ket{C}$).
Proof.
Let the set $\mathcal{O}$ contain $U_{\_}{C^{*}}$ for every possible subspace $C\subseteq\mathbb{F}^{n}_{\_}2$ such that $\dim(C)=n/2$. Let $(U_{\_}{C^{*}},U_{\_}{D^{*}})\in R$ if and only if $\dim(C\cap D)=n/2-1$. Then given $U_{\_}{C^{*}}\in\mathcal{O}$ and |$\ket{\eta}\in S^{\perp}_{\_}{C^{*}}$, let
|$\ket{\eta}=\sum_{\_}{x\in\{0,1\}^{n+k}\setminus C^{*}}\alpha_{\_}x\ket{x}.$
We have
$\mathop{\mathbb{E}}_{\_}{U_{\_}{D^{*}}\,:\,(U_{\_}{C^{*}},U_{\_}{D^{*}})\in R}\left[F(\ket{\eta},S_{\_}{D^{*}})^{2}\right]$ $=\mathop{\mathbb{E}}_{\_}{D\,:\,\dim(D)=n/2,\,\dim(C\cap D)=n/2-1}\left[\sum_{\_}{x\in D^{*}\setminus C^{*}}|\alpha_{\_}x|^{2}\right]$ (26)
$\leq\max_{\_}{x\in\{0,1\}^{n+k}\setminus C^{*}}\bigg{(}\Pr_{\_}{D\,:\,\dim(D)=n/2,\,\dim(C\cap D)=n/2-1}[x\in D^{*}]\bigg{)}$ (27)
$=\max_{\_}{x\in\{0,1\}^{n+k-1}\setminus C^{*}_{\_}{\mathcal{E}_{\_}X}}\bigg{(}\Pr_{\_}{D\,:\,\dim(D)=n/2,\,\dim(C\cap D)=n/2-1}[x\in D^{*}_{\_}{\mathcal{E}_{\_}X}]\bigg{)}$ (28)
$=\frac{|D^{*}_{\_}{\mathcal{E}_{\_}X}\setminus C^{*}_{\_}{\mathcal{E}_{\_}X}|}{|\{0,1\}^{n+k-1}\setminus C^{*}_{\_}{\mathcal{E}_{\_}X}|}$ (29)
$=\frac{|D^{*}_{\_}{\mathcal{E}_{\_}X}\setminus C^{*}_{\_}{\mathcal{E}_{\_}X}|}{|\mathcal{E}_{\_}X|2^{n}-|\mathcal{E}_{\_}X|2^{n/2}}$ (30)
|_$\leq\frac{|\mathcal{E}_{\_}X|2^{n/2}-2^{n/2-1}}{|\mathcal{E}_{\_}X|2^{n}-|\mathcal{E}_{\_}X|2^{n/2}}$ (31)
=|_$=\frac{|\mathcal{E}_{\_}X|-2^{-1}}{|\mathcal{E}_{\_}X|\left(2^{n/2}-1\right)}$ (32)
$\leq\frac{1}{2^{n/2}-1}$ (33)
$\leq\frac{1}{2^{n/2-1}}$ (34)
Equation 26 uses the definition of the fidelity. As done in [AC12], equation 27 uses the easy direction of the minimax theorem. Equation 28 uses the symmetry between $C^{*}_{\_}{\mathcal{E}_{\_}X}$ and $C^{*}_{\_}{\mathcal{E}_{\_}Z}$. Equation 30 uses equation 24 and the following facts: each coset has the same number of elements as the group; two cosets for the same group are either disjoint or identical; and that there is a coset for each error vector. Equation 31 uses the fact that $|D^{*}_{\_}{\mathcal{E}_{\_}X}\setminus C^{*}_{\_}{\mathcal{E}_{\_}X}|=|D^{*}_{\_}{\mathcal{E}_{\_}X}|-|D^{*}_{\_}{\mathcal{E}_{\_}X}\cap C^{*}_{\_}{\mathcal{E}_{\_}X}|$ and that the intersection of two cosets from $D$ and $C$ are either empty or a coset of $D\cap C$; while we do not know how many distinct coset intersections we have, we know that there is at least one such intersection: the trivial coset $C\cap D$ which has $2^{n/2-1}$ elements. The conclusion here is that $\varepsilon\coloneqq 2^{-n/2+1}$.
[…]
$\braket{C_{\_}{e,e^{\prime}}}{D_{\_}{t,t^{\prime}}}$ $=\bigg{(}\frac{1}{\sqrt{|C|}}\sum_{\_}{v\in C}(-1)^{v\cdot e^{\prime}}\bra{v+e},\frac{1}{\sqrt{|D|}}\sum_{\_}{z\in D}(-1)^{z\cdot t^{\prime}}\ket{z+t}\bigg{)}$
$=\frac{1}{2^{n/2}}\bigg{(}\sum_{\_}{v\in C}(-1)^{v\cdot e^{\prime}}\bra{v+e},\sum_{\_}{z\in D}(-1)^{z\cdot t^{\prime}}\ket{z+t}\bigg{)}$
$\leq\frac{1}{2^{n/2}}\sum_{\_}{v\in C}\sum_{\_}{z\in D}\braket{v+e}{z+t}$
Now suppose that the intersection of the two cosets, $C+e\cap D+t$ is non-empty. Then we can interpret the above sum of inner-products as the number of elements in a coset of $C\cap D$, which is equal to $|C\cap D|=2^{n/2-1}$. Hence,
$\braket{C_{\_}{e,e^{\prime}}}{D_{\_}{t,t^{\prime}}}\leq\frac{2^{n/2-1}}{2^{n/2}}=\frac{1}{2},$
which indeed implies $|\braket{f_{\_}C}{f_{\_}D}|\leq 1/4$.
Then by 4.3, with $c=1/2,d=1/4,$ and $\varepsilon=2^{-n/2+1}$,
$\Omega\left(\frac{c-d}{\sqrt{\varepsilon}}\right)=\Omega\left(2^{\left(n-2\right)/4}\right)=\Omega\left(2^{n/4}\right)$
The following corollary shows that, for the case where the counterfeiter succeeds almost perfectly, exponentially many queries are still needed.
Corollary 4.5 (Lower bound for small-error counterfeiting).
Given one copy of |$\ket{C}$, as well as oracle access to $U_{\_}{C^{*}}$, a counterfeiter needs $\Omega(2^{n/4})$ queries to prepare a state $\rho$ such that
$\Big{(}\bra{C_{\_}{e,e^{\prime}}}\otimes\bra{C_{\_}{t,t^{\prime}}}\Big{)}\rho\Big{(}\ket{C_{\_}{e,e^{\prime}}}\otimes\ket{C_{\_}{t,t^{\prime}}}\Big{)}\geq 1-\epsilon$
for small $\epsilon$ (say $\epsilon=0.0001$) and some $(e,e^{\prime}),(t,t^{\prime})\in\mathcal{E}$, for a worst case |$\ket{C}$. The probability that this state passes verification is
$\Pr[V^{\otimes 2}(\rho)=\text{accepts}]\geq 1-\epsilon.$
Proof.
Let $|\braket{C}{D}|=c$. If the counterfeiter succeeds, it must map |$\ket{C}$ to some state $\rho_{\_}C$ such that
$\Big{(}\bra{C_{\_}{\tilde{e},\tilde{e}^{\prime}}}\otimes\bra{C_{\_}{\tilde{t},\tilde{t}^{\prime}}}\Big{)}\rho_{\_}C\Big{(}\ket{C_{\_}{\tilde{e},\tilde{e}^{\prime}}}\otimes\ket{C_{\_}{\tilde{t},\tilde{t}^{\prime}}}\Big{)}\geq 1-\epsilon$
for some $(\tilde{e},\tilde{e}^{\prime}),(\tilde{t},\tilde{t}^{\prime})\in\mathcal{E}$. Likewise, it must map |$\ket{D}$ to some state $\rho_{\_}D$ such that
$\Big{(}\bra{C_{\_}{\tilde{s},\tilde{s}^{\prime}}}\otimes\bra{C_{\_}{\tilde{r},\tilde{r}^{\prime}}}\Big{)}\rho_{\_}D\Big{(}\ket{C_{\_}{\tilde{s},\tilde{s}^{\prime}}}\otimes\ket{C_{\_}{\tilde{r},\tilde{r}^{\prime}}}\Big{)}\geq 1-\epsilon$
for some $(\tilde{s},\tilde{s}^{\prime}),(\tilde{r},\tilde{r}^{\prime})\in\mathcal{E}$
Let |$\ket{f_{\_}C},\ket{f_{\_}D}$ be purifications of $\rho_{\_}C,\rho_{\_}D$ respectively. Then by 2.2
$|\braket{f_{\_}C}{f_{\_}D}|$ | $\leq F(\rho_{\_}C,\rho_{\_}D)$
$\leq\Bigg{|}\bra{C_{\_}{\tilde{e},\tilde{e}^{\prime}}}\otimes\bra{C_{\_}{\tilde{t},\tilde{t}^{\prime}}}\ket{D_{\_}{s,s^{\prime}}}\otimes\ket{D_{\_}{r,r^{\prime}}}\bigg{|}+2\epsilon^{1/4}$
$\leq c^{2}+2\epsilon^{1/4}$
where the upper bound is achieved by assuming that the above coset intersections are non-empty. Thus, with $d\coloneqq c^{2}+2\epsilon^{1/4}$ in 4.3,
$\Omega\bigg{(}\frac{c-d}{\sqrt{\varepsilon}}\bigg{)}=\Omega\left(\left(c-c^{2}-2\epsilon^{1/4}\right)\left(2^{\left(n-2\right)/4}\right)\right).$
Fixing $c=1/2$ gives $\Omega\left(2^{n/4}\right)$
The probability that such a state $\rho$ passes verification is
$\Pr[V^{\otimes 2}(\rho)=\text{accepts}]$ =_$=\sum_{\_}{(e,e^{\prime})^{\prime},(t,t^{\prime})\in\mathcal{E}}\bra{C_{\_}{e,e^{\prime}}}\otimes\bra{C_{\_}{t,t^{\prime}}}\rho\ket{C_{\_}{e,e^{\prime}}}\otimes\ket{C_{\_}{t,t^{\prime}}}$
$\geq\left(\bra{C_{\_}{\tilde{e},\tilde{e}^{\prime}}}\otimes\bra{C_{\_}{\tilde{t},\tilde{t}^{\prime}}}\Big{)}\rho\Big{(}\ket{C_{\_}{\tilde{e},\tilde{e}^{\prime}}}\otimes\ket{C_{\_}{\tilde{t},\tilde{t}^{\prime}}}\right)$
$\geq 1-\epsilon.$
Recall the implication of theorem 4.1: since a counterfeiter with small success probability can be amplified to one with high success probability, it suffices to prove security against the latter. We now formally use this theorem, and the previous result on counterfeiters that succeed with high probability (corollary 4.5), to provide a lower bound to the number of queries needed by a counterfeiter that has small success probability.
While the proof technique for this result is practically the same as the analogous result in [AC12], our result contains a noticeable difference. Since the space $S$ which our verification operator projects to is not 1-dimensional, the probability of a counterfeiter’s state being accepted can be notably higher in comparison to the noise-free setting, even if that state has a low fidelity with the tensor product of two tolerable noisy states. This difference marks the trade-off between security and noise-tolerance.
Corollary 4.6 (Lower bound for high-error counterfeiting).
Let $1/\varepsilon=o(2^{n/2})$. Given one copy of |$\ket{C}$, as well as oracle access to $U_{\_}{C^{*}}$, a counterfeiter needs $\Omega\left(\sqrt{\varepsilon}2^{n/4}\right)$ queries to prepare a state $\rho$ such that
$\max_{\_}{(e,e^{\prime}),(t,t^{\prime})\in\mathcal{E}}\Big{(}\bra{C_{\_}{e,e^{\prime}}}\otimes\bra{C_{\_}{t,t^{\prime}}}\Big{)}\rho\Big{(}\ket{C_{\_}{e,e^{\prime}}}\otimes\ket{C_{\_}{t,t^{\prime}}}\Big{)}\geq\varepsilon$
for a worst-case |$\ket{C}$. Furthermore, the greatest lower bound on the probability of accepting this state is
$\Pr[V^{\otimes 2}(\rho)=\text{accepts}]\geq|\mathcal{E}|^{2}\varepsilon.$
Proof.
Suppose we have a counterfeiter $\mathscr{C}$ that makes $o\left(\sqrt{\varepsilon}2^{n/4}\right)$ queries to $U_{\_}{C^{*}}$, and prepares a state $\sigma$ such that
$\max_{\_}{(e,e^{\prime}),(t,t^{\prime})\in\mathcal{E}}\Big{(}\bra{C_{\_}{e,e^{\prime}}}\otimes\bra{C_{\_}{t,t^{\prime}}}\Big{)}$ $\sigma\Big{(}\ket{C_{\_}{e,e^{\prime}}}\otimes\ket{C_{\_}{t,t^{\prime}}}\Big{)}\geq\varepsilon$
$\Longrightarrow F(\sigma,S^{\otimes 2})^{2}$ $\geq\varepsilon.$
Then
$\Pr[V^{\otimes 2}(\sigma)=\text{accepts}]=F(\sigma,S^{\otimes 2})^{2}$ $\geq\varepsilon.$
Let $\delta$ be small, say $\delta=0.00001$. Then by theorem 4.1, there exists an amplified counterfeiter $\mathscr{C}^{\prime}$ that makes
$\mathcal{O}\left(\frac{\log 1/\delta}{\sqrt{\varepsilon}\left(\sqrt{\varepsilon}+\delta^{2}\right)}\right)=\mathcal{O}\left(\frac{1}{\sqrt{\varepsilon}}\right)$
calls to $\mathscr{C}$ and $V_{\_}C$, and that prepares a state $\rho$ such that $F(\rho,S^{\otimes 2})\geq 1-\delta$ which implies
$\Pr[V^{\otimes 2}(\rho)=\text{accepts}]\geq(1-\delta)^{2}.$
Counting the $o\left(\sqrt{\varepsilon}2^{n/4}\right)$ queries from each use of $\mathscr{C}$, and $\mathcal{O}(1)$ queries from each use of $V_{\_}C$, the total number of queries that $\mathscr{C}^{\prime}$ makes to $U_{\_}{C^{*}}$ is
$\left[o\left(\sqrt{\varepsilon}2^{n/4}\right)+\mathcal{O}(1)\right]\cdot\mathcal{O}\left(\frac{1}{\sqrt{\varepsilon}}\right)=o\left(2^{n/4}\right).$
But this contradicts corollary 4.5.
Now suppose a counterfeiter uses $\Omega\left(\sqrt{\varepsilon}2^{n/4}\right)$ queries to prepare a state $\rho$ such that
$\max_{\_}{(e,e^{\prime}),(t,t^{\prime})\in\mathcal{E}}\Big{(}\bra{C_{\_}{e,e^{\prime}}}\otimes\bra{C_{\_}{t,t^{\prime}}}\Big{)}\rho\Big{(}\ket{C_{\_}{e,e^{\prime}}}\otimes\ket{C_{\_}{t,t^{\prime}}}\Big{)}\geq\varepsilon.$
If $\mathscr{C}$ submits this state for verification, their probability of acceptance will be higher if
$\Big{(}\bra{C_{\_}{e,e^{\prime}}}\otimes\bra{C_{\_}{t,t^{\prime}}}\Big{)}\rho\Big{(}\ket{C_{\_}{e,e^{\prime}}}\otimes\ket{C_{\_}{t,t^{\prime}}}\Big{)}>0,$
for all $(e,e^{\prime}),(t,t^{\prime})\in\mathcal{E}$. The best scenario for $\mathscr{C}$ is that each term obtains the maximum value, in which case we have
$\Pr[V^{\otimes 2}(\rho)=\text{accepts}]$ =_$=\sum_{\_}{(e,e^{\prime}),(t,t^{\prime})\in\mathcal{E}}\bra{C_{\_}{e,e^{\prime}}}\otimes\bra{C_{\_}{t,t^{\prime}}}\rho\ket{C_{\_}{e,e^{\prime}}}\otimes\ket{C_{\_}{t,t^{\prime}}}$
$=|\mathcal{E}|^{2}\cdot\max_{\_}{(e,e^{\prime}),(t,t^{\prime})\in\mathcal{E}}\Big{(}\bra{C_{\_}{e,e^{\prime}}}\otimes\bra{C_{\_}{t,t^{\prime}}}\Big{)}\rho\Big{(}\ket{C_{\_}{e,e^{\prime}}}\otimes\ket{C_{\_}{t,t^{\prime}}}\Big{)}$
$\geq|\mathcal{E}|^{2}\varepsilon.$
The following theorem shows that it suffices for a bank to generate uniformly random banknotes. Indeed, as the proof shows, if a counterfeiter is able to clone a uniformly random banknote, then they can clone any banknote. Note that in the noise-free case of [AC12], the proof of this theorem used an invertible linear map $f$; in our case, we additionally require $f$ to be an isometry (i.e., a distance-preserving map) to ensure that the noise-tolerance is, in a sense, “preserved” under the mapping $f$.
Theorem 4.7 (Lower bound for average-case counterfeiting).
Let $C$ be a uniformly random element of $\mathcal{W}$. Then given one copy of |$\ket{C}$, as well as oracle access to $U_{\_}{C^{*}}$, a counterfeiter $\mathscr{C}$ needs $\Omega(\sqrt{\varepsilon}2^{n/4})$ queries to prepare a $2n$-qubit state $\rho$ that $V^{\otimes 2}_{\_}C$ accepts with probability at least $|\mathcal{E}|^{2}\varepsilon$, for all $1/\varepsilon=o(2^{n/2})$. Here the probability is taken over the choice $C$, as well as the behaviour of $\mathscr{C}$ and $V^{\otimes 2}_{\_}C$.
Proof.
Suppose we had a counterfeiter $\mathscr{C}$ that violated the above. Using $\mathscr{C}$ as a black box, we construct a new counterfeiter $\mathscr{C}^{\prime}$ that violates corollary 4.6.
Given a (deterministically-chosen) banknote |$\ket{C}$ and oracle access to $U_{\_}{C^{*}}$, first choose an invertible linear isometry $f:\mathbb{F}^{n}_{\_}2\rightarrow\mathbb{F}^{n}_{\_}2$ uniformly at random. Then $f(C)$ is a uniformly random element of $\mathcal{W}$ by 3.6. Furthermore, the state |$\ket{C}$ can be transformed into |$\ket{f(C)}$ straightforwardly.
Using the oracle $U_{\_}{C^{*}}$ and the map $f$, we can simulate the corresponding oracle for the money state |$\ket{f(C)}$. Let $U_{\_}f$ be the unitary that acts as $U_{\_}f\ket{x}=\ket{f(x)}$. Then the corresponding oracle is given by $(\mathds{1}^{k}\otimes U_{\_}f)U_{\_}{C^{*}}(\mathds{1}^{k}\otimes U^{\dagger}_{\_}f)$. To see that this works, observe that,
$x\in C_{\_}{\mathcal{E}_{\_}X}$ $\Leftrightarrow x\in C+e,\text{ for some }e\in\mathcal{E}_{\_}X$
$\Leftrightarrow f(x)\in f(C)+f(e),\text{ where }f(e)\in\mathcal{E}_{\_}X$
$\Leftrightarrow f(x)\in f(C)_{\_}{\mathcal{E}_{\_}X}$
where $f(e)\in\mathcal{E}_{\_}X$ follows because $f$ is an isometry and thus preserves the weight of $e$.
So, by using $\mathscr{C}$ for $n/2$-dimensional states that are drawn uniformly randomly from $\mathcal{W}$, the adversary $\mathscr{C}^{\prime}$ can produce a state $\rho_{\_}f$ that $V^{\otimes 2}_{\_}{f(C)}$ accepts with probability at least $|\mathcal{E}|^{2}\varepsilon$. Applying $f^{-1}$ to both registers of $\rho_{\_}f$, $\mathscr{C}^{\prime}$ can obtain a state $\rho$ that $V^{\otimes 2}_{\_}C$ accepts with probability at least $|\mathcal{E}|^{2}\varepsilon$, thereby contradicting corollary 4.6. ∎
We now rephrase theorem 4.7 in terms of the oracle pair $(U_{\_}{C_{\_}{\mathcal{E}_{\_}X}},U_{\_}{C_{\_}{\mathcal{E}_{\_}Z}})$ as opposed to the single oracle $U_{\_}{C^{*}}$. In the noise-free case of [AC12], this is immediate as their single oracle $U_{\_}{A^{*}}$ exactly corresponds to the pair $(U_{\_}A,U_{\_}{A^{\perp}})$. In our case, we recall (see section 4.1.2) that the query complexity for the pair $(U_{\_}{C_{\_}{\mathcal{E}_{\_}X}},U_{\_}{C_{\_}{\mathcal{E}_{\_}Z}})$ is the query complexity for $U_{\_}{C^{*}}$ divided by $2|\mathcal{E}_{\_}X|$. Under this correspondence, we get the following corollary of theorem 4.7.
Corollary 4.8.
Let $C$ be a uniformly random element of $\mathcal{W}$. Then given one copy of |$\ket{C}$, as well as oracle access to $(U_{\_}{C_{\_}{\mathcal{E}_{\_}X}},U_{\_}{C_{\_}{\mathcal{E}_{\_}Z}})$, a counterfeiter $\mathscr{C}$ needs $\Omega(\sqrt{\varepsilon}2^{n/4}/|\mathcal{E}_{\_}X|)$ queries to prepare a $2n$-qubit state $\rho$ that $V^{\otimes 2}_{\_}C$ accepts with probability at least $|\mathcal{E}|^{2}\varepsilon$, for all $1/\varepsilon=o(2^{n/2})$. Here the probability is taken over the choice $C$, as well as the behaviour of $\mathscr{C}$ and $V^{\otimes 2}_{\_}C$.
In the noise-free case, corollary 4.8 implies that $\varepsilon$ is $1/\textsf{exp}(n)$, which constitutes the soundness error in this case. In the noisy case, we need $|\mathcal{E}|^{2}$ to be constant in $n$ to get the same soundness error, i.e., for $|\mathcal{E}|^{2}\varepsilon$ to be $1/\textsf{exp}(n)$. More errors can be tolerated at the cost of increasing the soundness error, though we remark that it is not possible to tolerate too many errors. Recall from equation 11 that the size of |_$|\mathcal{E}_{\_}q|$ is given by
|_$|\mathcal{E}_{\_}q|=|\mathcal{E}_{\_}X|\cdot|\mathcal{E}_{\_}Z|=\left(\sum^{q}_{\_}{j=0}\binom{n}{j}\right)^{2}.$
When $q$ is large (close to $n$), the sum in the parentheses approaches exponential. This trade-off between noise-tolerance and soundness is visualized in figure 2. We see that if we fix a desirable value for the soundness error, then increasing $q$ necessitates an increase of $n$ to maintain the same soundness error. In addition to increasing the soundness error, corollary 4.8 shows that tolerating more errors has the effect of reducing the number of queries needed by an adversary.
Figure 2: A plot of the soundness error |_$|\mathcal{E}_{\_}q|^{2}\varepsilon$ for various choices of $q$ (the number of tolerated arbitrary errors) against the number of qubits $n$. Here, we take $\varepsilon=2^{-n/2}$.
We now return to proving security for the noisy-projective mini-scheme $\mathcal{M}$ specified in section 3.5. To do this, we use the same proof technique as the analogous result in [AC12]; the notable difference here is that we must account for the fact that the formal specification of our scheme uses conjugate coding states, though this does not affect security.
Theorem 4.9 (Security of the noisy-projective mini-scheme).
The noisy-projective mini-scheme $\mathcal{M}=(\textsf{Bank}_{\_}\mathcal{M},\textsf{Ver}_{\_}\mathcal{M})$, which is defined relative to the classical oracle $U$, has perfect completeness and $|\mathcal{E}|^{2}/\textsf{exp}(n)$ soundness error.
Proof.
Perfect completeness follows from how $\mathcal{M}$ is defined (see section 3.5) and from 3.5.
[…]
The goal is to show that any successful attack on $\mathcal{M}$ can be converted into successful counterfeiting of |$\ket{C}$ given oracle access to $(U_{\_}{C_{\_}{\mathcal{E}_{\_}X},r},U_{\_}{C_{\_}{\mathcal{E}_{\_}Z},r})$. This would then contradict corollary 4.8, thus proving the result. This can be achieved by observing that an adversary trying to counterfeit |$\ket{C}$, given $(U_{\_}{C_{\_}{\mathcal{E}_{\_}X},r},U_{\_}{C_{\_}{\mathcal{E}_{\_}Z},r})$, can essentially simulate the oracles $\mathcal{G},\mathcal{Z},\mathcal{T}_{\_}{\text{primal}},\mathcal{T}_{\_}{\text{dual}}$.
[…]
Thus, with these simulated oracles, any successful attack on $\mathcal{M}$ can indeed be converted into successful counterfeiting of |$\ket{C}$ with oracle access to only $(U_{\_}{C_{\_}{\mathcal{E}_{\_}X},r},U_{\_}{C_{\_}{\mathcal{E}_{\_}Z},r})$. This contradicts corollary 4.8. ∎
Remark 4.10.
The scheme we have presented tolerates noise but does not correct it. Consequently, a valid banknote will eventually accumulate too many errors to be accepted. It is straightforward to adapt the subset approach to correct tolerable noise. Instead of checking bit-flip errors with a single oracle $U_{\_}{C_{\_}{\mathcal{E}_{\_}X}}$, we would use a collection of oracles $\{U_{\_}{C+e}\}_{\_}{e\in\mathcal{E}_{\_}X}$ where $U_{\_}{C+e}$ checks for membership in the coset $C+e$, and similarly for phase-flip errors. Verification would involve querying these oracles sequentially; by doing this, we can determine if the money state is in one of the cosets (in which case we accept it) and, based off which coset it belongs to, we learn which error has occurred (which allows for correction). The security proof would proceed in a similar manner. In section 4.1.2, the construction of $C^{*}$ would remain the same, but now a single query to $U_{\_}{C+e}$ corresponds to a single query to $U_{\_}{C^{*}}$, and so the query complexity in corollary 4.8 would be $\Omega(\sqrt{\varepsilon}2^{n/4})$. The downside of this approach is that verification now involves the oracles ($(\{U_{\_}{C+e}\}_{\_}{e\in\mathcal{E}_{\_}X},\{U_{\_}{C+e}\}_{\_}{e\in\mathcal{E}_{\_}Z})$, whereas the scheme we presented for tolerating noise only involved $(U_{\_}{C_{\_}{\mathcal{E}_{\_}X}},U_{\_}{C_{\_}{\mathcal{E}_{\_}Z}})$, and so this scheme for error correction could involve many additional queries as part of the verification algorithm.
5 Future directions
In section 3.4, we presented two approaches to verification: the subset approach and the subspace approach. Is one preferable over the other?
Both approaches achieve the same completeness and soundness error, but is one easier to realize in the plain model? Towards answering this question, we note that by following the approach Zhandry took in [Zha19] to instantiate the oracle scheme of [AC12] with iO, it seems likely that the subspace approach could be similarly instantiated with iO. Is it any more difficult to use iO to instantiate the classical oracle that checks membership in subsets?
How does adding a layer of error correction compare to our approach?
In section 3.3.2, we briefly discussed how one could take a different approach to noise-tolerance by applying a layer of error correction on top of the [AC12] oracle scheme. This approach has no known security proof, though we have identifed the following obstacles: translating any necessary computations to versions that operate on encoded states, and determining if/how the original security proof changes when the adversary holds an encoded state instead of a subspace state. Given a realization of this approach, it would be interesting to make a comparison with our scheme. Notably, a layer of error correction typically requires many additional qubits (i.e., large overhead), so one may ask questions like: for a fixed $n$, how does the noise-tolerance and soundness error compare?
Is it possible to obtain noise-tolerance for other public-key quantum money schemes?
In section 1.3, we mentioned some of the other public-key quantum money schemes that rely on hardness assumptions instead of oracles. Is it possible to make any of these schemes noise-tolerant?
Acknowledgments
We would like to thank Anne Broadbent for advice and related discussions. We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), the US Air Force Office of Scientific Research under award number FA9550-20-1-0375, and the University of Ottawa’s Research Chairs program.